### v2

# A central limit theorem for partitions involving generalised divisor functions

Abstract.
We define an $f$-restricted partition $p_{f}(n,k)$ of fixed length $k$ given by the bivariate generating series
$Q_{f}(z,u)\coloneqq 1+\sum_{n=1}^{\infty}\sum_{k=1}^{\infty}p_{f}(n,k)u^{k}z^{n}=\prod_{k=1}^{\infty}(1+uz^{k})^{\Delta_{f}(k)},$
where $\Delta_{f}(n)=f(n+1)-f(n)$. In this article, we establish a central limit theorem for the number of summands in such partitions when $f(n)=\sigma_{r}(n)$ denotes the generalised divisor function, defined as $\sigma_{r}(n)=\sum_{d|n}d^{r}$ for integer $r\geq 2$. This can be considered as a generalisation of the work of Lipnik, Madritsch, and Tichy, who previously studied this problem for $f(n)=\lfloor{n}^{\alpha}\rfloor$ with $0<\alpha<1$. A key element of our proof relies on the analytic behaviour of the Dirichlet series
$\sum_{n=1}^{\infty}\frac{\sigma_{r}(n+1)}{n^{s}},$
for $\mathrm{Re}(s)>1$. We study this problem employing the identity involving the Ramanujan sum. Furthermore, we analyse the Euler product arising from the above Dirichlet series by adopting the argument of Alkan, Ledoan and Zaharescu.
Key words and phrases:
Central limit theorem, partition function, Dirichlet series with a shifted coefficient, Euler product, saddle-point method, converse mapping
2020 Mathematics Subject Classification:
11P82, 05A17, 60F05, 11M41
To Bruce Berndt, on the occasion of his 86th birthday
1. Introduction
An arithmetic function $f:\mathbb{N}\to\mathbb{R}_{>0}$ is called multiplicative if $f(mn)=f(m)f(n)$ whenever $(m,n)=1$. An integer partition $p_{f}(n)$ of a positive integer $n$ into the function $f$ counts the number of ways to write $n$ as follows
$n=\lfloor{f(n_{1})}\rfloor+\lfloor{f(n_{2})}\rfloor+\cdots+\lfloor{f(n_{r})}\rfloor,$ (1.1)
where $f(n_{j})$ is a real positive valued multiplicative function and $n_{1}\leq n_{2}\leq\cdots\leq n_{r}$. It is a well-known fact that one can write the above expression with the generating series
${Q_{f}}(z)=1+\sum_{n=1}^{\infty}{{p_{f}}(n){z^{n}}}=\prod_{n\in\mathbb{N}^{*}}(1-z^{\lfloor{f(n)}\rfloor})^{-1}\quad(|z|<1).$ (1.2)
The asymptotic formula for integer partitions, that is, when $f(n)=n$ for all $n\in\mathbb{N}$ in (1.2), was first established by Hardy and Ramanujan [15]. Over the past century, the study of partition asymptotics has attracted the attention of numerous mathematicians; see, for example, [8, 10, 11, 32]. The most general framework in this context was studied by Debruyne and Tenenbaum [12].
Let $p_{f}(n)$ and $p_{f}(n,r)$ denote, respectively, the “restricted” partition of $n$ and and the “restricted” partition of $n$ with fixed length $r$ as defined in (1.1), where $n_{1}\leq n_{2}\leq\ldots\leq n_{r}$. Li and Chen [20, Theorem 1.1] showed an alternative approach to analysing the generating series (1.2), which is through the study of
$Q_{f}(z)=\sum_{n=0}^{\infty}p_{f}(n)z^{n}=\prod_{n=1}^{\infty}(1-z^{n})^{-\Delta_{f}(n)}\quad(|z|<1).$ (1.3)
where
$\Delta_{f}(n)=\#\{m:m\in\mathbb{N}\;\text{and}\;\lfloor{f(m)}\rfloor=n\}.$ (1.4)
A related problem was studied by Balasubramanian and Luca [3], who investigated the number of solutions to $|\{m:m,n\in\mathbb{N}\quad\text{and}\quad m=\mathcal{F}(n)\}|$, where $\mathcal{F}(n)$ denotes the number of unordered factorisation as the products of integer factors greater than 1, by examining the partitions of the form $n=\lfloor{\sqrt{n_{1}}}\rfloor+\cdots+\lfloor{\sqrt{n_{r}}}\rfloor$. Later, this result was thoroughly investigated by Li and Chen [21].
Our main objective of this paper is to establish a central limit theorem for the random variable $\varpi_{f,n}$, which counts the number of summands in a random $f$-restricted partition of $n$ (as defined in Section 2) with respect to the multiplicative function $f$. This concept was first introduced by Erdős and Lehner [13], where they studied the distribution of ratio $p(n,r)/p(n)$, with $r=(2\pi^{2}/3)^{-\tfrac{1}{2}}\sqrt{n}\log n+n\sqrt{n}$ as a function of $n$ and $p(n)$ denoting number of integer partitions of $n$. In this context, it is noteworthy that the weighted moments of integer partitions were studied by Richmond [29, 30]. Let $p_{A}(n,r)$ denotes the partition of $n$ into $r$ summands from a given set $A$. The $k$-th moment is then defined by
$t^{k}_{A}(n)=\sum_{r=1}^{n}r^{k}p_{A}(n,r),$ (1.5)
for arbitrary fixed $k$ as $n\to\infty$. Furthermore, Richmond remarked that the study of these weighted moments supports the work of Erdős and Lehner.
In a recent study, Lipnik, Madritsch, and Tichy [22] established a central limit theorem for the partition function $p_{f}(n)$ with $f(n)=\lfloor{n^{\alpha}}\rfloor$ for $0<\alpha<1$. For our main result, we have considered $f$ to be the generalised divisor function $\sigma_{r}(n)$ for integers $r\geq 2$. However, our approach is applicable for a wider class of multiplicative functions, which we explore in detail in later sections.
The proof presented in [22] used a probabilistic approach and Curtiss’ theorem to obtain the asymptotics of the moment generating function. Furthermore, the application of converse mapping to study the analytic behaviour of the associated Dirichlet series makes their proof simple and elegant. This also allows us to choose a complex weight function in the bivariate generating series and deal with a challenging associated Dirichlet series, as the remaining components of the proof are comparatively less involved than those encountered in prior works on this problem.
Studying the limiting distribution of $p_{f}(n)$ for $f$ being any multiplicative function has a different flavour compared to previous works. Our analysis relies on additive combinatorics, due to the involvement of the gap properties $(|f(n+1)-f(n)|)_{n}$ of the multiplicative function $f$. While we establish a central limit theorem, deriving an asymptotic formula for (1.5) in the setting of $f$-restricted partitions would present additional difficulties, largely due to the behaviour of the gap function $(|f(n+1)-f(n)|)_{n}$. However, it appears plausible that an upper bound could be obtained in future work.
Throughout this paper $f\sim g$ means ${f(x)}/{g(x)}\to 1$ as $x\to\infty$. Additionally, $f=o(g)$ and $f=O(g)$ denote $\left|{f(x)}/{g(x)}\right|\to 0$ and $|f(x)|\leq C|g(x)|$ as $x\to\infty$, respectively, for a suitable constant $C>0$.
2. The Main Results
Let
$c_{m}(n)=\sum_{\begin{subarray}{c}b(\bmod{m})\\ (b,m)=1\end{subarray}}e\left(\frac{bn}{m}\right),$ (2.1)
denote the Ramanujan sum, and the $r$-th divisor function is given by [^1] In Section 4.1, we provide a detailed rationale for selecting this particular function within the context of our analysis.
$\sigma_{r}(n)=\sum_{d|n}d^{r}.$
Define
$\Delta_{f}(n)\coloneqq f(n+1)-f(n).$ (2.2)
Let $f$ be a positive integer valued multiplicative function. We define a $f$-restricted partition $p_{f}(n)$ such that the fixed-length partition $p_{f}(n,k)$ is associated with the bivariate generating series
$Q_{f}(z,u)\coloneqq 1+\sum_{n=1}^{\infty}\sum_{k=1}^{\infty}p_{f}(n,k)u^{k}z^{n}=\prod_{k=1}^{\infty}(1+uz^{k})^{\Delta_{f}(k)}.$ (2.3)
Let $\Pi(n)$ denote the collection of $f$-restricted partitions and let $p_{f}(n)=|\Pi(n)|$ and $p_{f}(n,k)=|\Pi(n,k)|$ denote the cardinalities of the sets $\Pi(n)$ and $\Pi(n,k)$, respectively. We define the random variable $\varpi_{f,n}$, which counts the number of summands in a random $f$-restricted partition. The probability distribution of $\varpi_{f,n}$ is given by
$\mathbb{P}(\varpi_{f,n}=k)=\frac{p_{f}(n,k)}{p_{f}(n)}.$
We derive the central limit theorem for the random variable $\varpi_{f,n}$ by examining the bivariate generating function $Q_{f}(z,u)$, which is defined in (2.3).
In the subsequent lemma, we establish the connection between the generating series $Q_{f}(z,u)$ and the random variable $\varpi_{f,n}$.
Lemma 2.1.
Consider the bivariate generating series $Q_{f}(z,u)$ as defined in (2.3). The following relation holds
$Q_{f}(z,u)=\prod_{k=1}^{\infty}(1+uz^{k})^{\Delta_{f}(k)}=1+\sum_{n=1}^{\infty}p_{f}(n)\mathbb{E}(u^{\varpi_{f,n}})z^{n}.$
Proof.
The first relation follows directly from (2.3). Furthermore rearranging the terms in (2.3) yields
$Q_{f}(z,u)=$ $1+\sum_{n=1}^{\infty}\sum_{k=1}^{\infty}p_{f}(n,k)u^{k}z^{n}$
$=$ $1+\sum_{n=1}^{\infty}p_{f}(n)z^{n}\sum_{k=1}^{\infty}\frac{p_{f}(n,k)}{p_{f}(n)}u^{k}$
$=$ $1+\sum_{n=1}^{\infty}p_{f}(n)\mathbb{E}(u^{\varpi_{f,n}})z^{n},$
thus concluding the lemma. ∎
Prior to stating our main result, we note that Lipnik et al. [22] considered the function $\Delta_{f}(n)=\lfloor{(n+1)^{\alpha}}\rfloor-\lfloor{n^{\alpha}}\rfloor$ for any real $0<\alpha<1$. We replace their weight function $\Delta_{f}(n)$ with $\sigma_{r}(n+1)-\sigma_{r}(n)$. Furthermore, an analogous result may be obtained using our method for any positive integer valued multiplicative function that satisfies an identity involving the Ramanujan sum.
Theorem 2.2.
Let $f(n)=\sigma_{r}(n)$ denote the $r$-th divisor function for fixed integer $r\geq 2$. Furthermore, let $\varpi_{f,n}$ be the random variable counting the number of summands in a random $f$-restricted partition of $n$ as defined in (2.3). Then, as $n\to\infty$, $\varpi_{f,n}$ is asymptotically normally distributed with mean $\mu_{f,n}$ and variance $\nu^{2}_{f,n}$ i.e.,
$\mathbb{P}\left(\frac{\varpi_{f,n}-\mu_{f,n}}{\nu_{f,n}}<x\right)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{x}e^{-{t^{2}}/{2}}dt+o(1),$
uniformly for all $x$ as $n\to\infty$. The mean $\mu_{f,n}$ and the variance $\nu^{2}_{f,n}$ are given by
$\mu_{f,n}=\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)}{e^{\eta k}+1}=O(C_{\mu}(r)n^{\frac{r+1}{r+2}})$ (2.4)
and
$\nu^{2}_{f,n}=\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)e^{\eta k}}{(e^{\eta k}+1)^{2}}-\frac{\left(\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)ke^{\eta k}}{(e^{\eta k}+1)^{2}}\right)^{2}}{\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)k^{2}e^{\eta k}}{(e^{\eta k}+1)^{2}}}=O(C_{\sigma}(r)n^{\frac{r+1}{r+2}}),$ (2.5)
where the constants $C_{\mu}(r)$ and $C_{\sigma}(r)$ are defined in (5.3) and (5.1), respectively, and $\eta$ is the implicit solution of
$n=\sum_{k\geq 1}\frac{k}{e^{\eta k}+1}.$
Additionally, the tail of the distribution is given by
$\mathbb{P}\left(\frac{\varpi_{f,n}-\mu_{f,n}}{\nu_{f,n}}\geq x\right)\leq\begin{cases}e^{-x^{2}/2}\left(1+O((\log n)^{-3})\right)&\text{if $0\leq x\leq n^{\frac{r+1}{6(r+2)}}/\log n$},\\ e^{-n^{\frac{r+1}{6(r+2)}}x/2}\left(1+O((\log n)^{-3})\right)&\text{if $x\geq n^{\frac{r+1}{6(r+2)}}/\log n$},\end{cases}$
and an analogous statement holds for $\mathbb{P}\left(\frac{\varpi_{f,n}-\mu_{f,n}}{\nu_{f,n}}\leq-x\right)$.
2.1. Initial Setup
In this section, we provide an outline of the proof of Theorem 2.2. We adopt the saddle-point approach from [22, §3] for the initial construction. However, our analysis necessitates a more intricate argument concerning the distribution of multiplicative functions due to the presence of the Dirichlet series with a shifted coefficient.
We begin by observing that, to establish the central limit theorem for the random variable $\varpi_{f,n}$, it is necessary to show that the normalised moment generating function $\mathbb{E}\left(e^{(\varpi_{f,n}-\mu_{f,n})t/\nu_{f,n}}\right)$ converges to $e^{t^{2}/2}$ as $t\to\infty$. Finally, we estimate the tail of the distribution using the Chernoff bound.
Since $\mu_{f,n}/\nu_{f,n}$ is a constant, we can write
$\mathbb{E}\left(e^{(\varpi_{f,n}-\mu_{f,n})t/\nu_{f,n}}\right)=e^{-\mu_{f,n}/\nu_{f,n}}\mathbb{E}\left(e^{\varpi_{f,n}t/\nu_{f,n}}\right).$
To achieve our goal we need to obtain the coefficient of $z^{n}$ in the bivariate generating series $Q(z,u)$. By Lemma 2.1, the probability generating function is expressed as
$\mathbb{E}(u^{\varpi_{f,n}})=\frac{[z^{n}]Q_{f}(z,u)}{[z^{n}]Q_{f}(z,1)}.$
Applying Cauchy integral formula we have
$\mathcal{Q}_{f,n}(u)\coloneqq[z^{n}]Q_{f}(z,u)=\frac{1}{2\pi i}\oint_{|z|=e^{-\tau}}\frac{Q_{f}(z,u)}{z^{n+1}}dz.$
A standard transformation gives, for $\tau>0$,
$\mathcal{Q}_{f,n}(u)=\frac{e^{n\tau}}{2\pi}\int_{-\pi}^{\pi}\exp\left(in\theta+F(\tau+i\theta,u)\right)d\theta,$ (2.6)
where
$F(\gamma,u)\coloneqq\log Q_{f}(e^{-\gamma},u)=\log\left(\prod_{k=1}^{\infty}(1+ue^{-\gamma k})^{\Delta_{f}(k)}\right).$ (2.7)
To tackle the integral in (2.6), we utilise the saddle-point technique, which is often referred to as the method of steepest descent. It can be viewed as a coarse version of the circle method. Accordingly, the primary application of this method is to derive estimates for integrals of the form
$\int_{-\pi}^{\pi}e^{\Phi(\tau+i\theta)}d\theta,$
for a suitable function $\Phi$. We will now split the integral into two parts. Choose $\theta_{n}>0$, then we write
$\int_{-\pi}^{\pi}e^{\Phi(\tau+i\theta)}d\theta=\int_{|\theta|\leq\theta_{n}}e^{\Phi(\tau+i\theta)}d\theta+\int_{\theta_{n}<|\theta|\leq\pi}e^{\Phi(\tau+i\theta)}d\theta.$ (2.8)
The first part will compute the contribution near the positive real axis, or, equivalently, around $\theta=0$. This will give us the main term, whereas the contribution from the real line i.e, the estimate arising from $e^{\Phi(\tau+i\theta)-\mathrm{Re}(\Phi(\tau))}$ will give us the error term. Using the Taylor expansion up to the third order, the function $\Phi$ can be expressed as
$\Phi(\tau+i\theta)=\Phi(\tau)+i\theta\Phi^{\prime}(\tau)-\frac{\theta^{2}}{2}\Phi^{\prime\prime}(\tau)+O\left(\sup_{\theta}\left|\theta^{3}\Phi^{\prime\prime\prime}(\tau+i\theta)\right|\right).$ (2.9)
We now choose $\tau$ such that the first-order derivative $\Phi^{\prime}(\tau)$ vanishes. Thus, the remaining term of the integral in (2.8) becomes
$\int_{|\theta|<\theta_{n}}e^{\Phi(\tau+i\theta)}d\theta=e^{\Phi(\tau)}\int_{|\theta|<\theta_{n}}e^{-\tfrac{\theta^{2}}{2}\Phi^{\prime\prime}(\tau)}\left(1+O\left(\sup_{\theta}\left|\theta^{3}\Phi^{\prime\prime\prime}(\tau+i\theta)\right|\right)\right)d\theta.$
We need to analyse the integrand with the terms involving $\Phi^{\prime\prime}(\tau)$ and $\Phi^{\prime\prime\prime}(\tau)$ in order to show that the aforementioned transformation and estimates are valid. For a detailed explanation, we refer to [22, §3]. Recall the Mellin transform of a function $h(t)$, defined by
$h^{*}(s)\coloneqq\mathcal{M}[h;s]=\int_{0}^{\infty}h(t)t^{s-1}dt,$
which serves as one of the key ingredients in our analysis. Of particular importance is the rescaling rule (see, for instance, [14, Theorem 1]), given by
$\mathcal{M}\left[\sum_{k}\lambda_{k}h(\mu_{k}x);s\right]=h^{*}(s)\sum_{k}\frac{\lambda_{k}}{\mu_{k}^{s}}.$
Let $\delta>0$ and set $\delta\leq u\leq\delta^{-1}$. By “uniformly in uu”, we mean “uniformly as $\delta\leq u\leq\delta^{-1}$” throughout this paper. Thus, the Mellin transform for the polynomial with respect to $\gamma$ yields
$\mathcal{M}\left[\sum_{k}\Delta_{f}(k)\log(1+ue^{-\gamma k});s\right]=\mathcal{D}_{f}(\Delta;s)\mathcal{I}(s,u),$
where
$\mathcal{D}_{f}(\Delta;s)\coloneqq\sum_{n=1}^{\infty}\frac{\Delta_{f}(n)}{n^{s}},$ (2.10)
and
$\mathcal{I}(s)=\int_{0}^{\infty}(1+ue^{-\gamma})\gamma^{s-1}d\gamma,$
represents the Mellin transform of $\gamma\to(1+ue^{-\gamma})$.
The principal advantage of the Mellin transform lies not in the transform itself, but in the inverse mapping, which allows us to examine the singularities of the transformed function to obtain the asymptotic expansion11footnotetext: We adopt the term ‘converse mapping’ in Theorem 2.3 from [22], where it is used in the context of the inverse Mellin transform.. This step is essential in our case, as we will be working with a Dirichlet series with a shifted multiplicative coefficient.
Theorem 2.3 (Converse Mapping [14]).
Let $f(x)$ be a continuous function on the interval $(0,+\infty)$, and suppose that its Mellin transform $f^{*}(s)$ has a non-empty fundamental strip $\langle\alpha,\beta\rangle$. Assume that $f^{*}(s)$ admits a meromorphic continuation to the larger strip $\langle\gamma,\beta\rangle$, for some $\gamma<\alpha$, with only finitely many poles in that region, and is analytic on the line $\mathrm{Re}(s)=\gamma$. Suppose further that there exists a real number $\eta\in(\alpha,\beta)$ such that
$f^{*}(s)=O(|s|^{-r})$ (2.11)
as $|s|\to\infty$ in $\gamma\leq\mathrm{Re}(s)\leq\eta$, for some $r>1$. If, in addition, $f^{*}(s)$ admits a singular expansion of the form
$f^{*}(s)\asymp\sum_{(\xi,k)\in A}\frac{d_{\xi,k}}{(s-\xi)^{k}}$ (2.12)
for $s\in\langle\gamma,\alpha\rangle$, then the function admits the following asymptotic expansion as $x\to 0^{+}$:
$f(x)=\sum_{(\xi,k)\in A}d_{\xi,k}\left(\frac{(-1)^{k-1}}{(k-1)!}x^{-\xi}(\log x)^{k-1}\right)+O(x^{-\gamma}).$
Therefore, in our case, we must consider the singularities of the Dirichlet series $\mathcal{D}_{f}(\Delta;s)$ and $\mathcal{I}(s,u)$. Thus, we are left with the expression
$\mathcal{I}(s,u)=\int_{0}^{\infty}(1+ue^{-\gamma})\gamma^{s-1}d\gamma=\int_{0}^{\infty}\sum_{j=1}^{\infty}\frac{(-u)^{j}}{j}e^{-j\gamma}\gamma^{s-1}d\gamma=\mathrm{Li}_{s+1}(-u)\Gamma(s),$
where $\mathrm{Li}_{s}(u)$ denotes the polylogarithm, and $\Gamma(s)$ represents the Gamma function. Now, we verify that conditions (2.11) and (2.12) hold for $\mathrm{Li}_{s}(u)$ and $\Gamma(s)$. An application of Stirling formula yields (see for instance [17, §5 (A.4)])
$|\Gamma(\sigma+it)|=\sqrt{2\pi}|t|^{\sigma-\tfrac{1}{2}}e^{-\pi|t|/2}(1+O_{a,b}(1/t)),$
where $a\leq\sigma\leq b$ and $|t|\geq 1$. For the polylogarithm function, we adopt the approach outlined in [22]. Let $w=-\log z$, and define
$\Lambda(w)\coloneqq\mathrm{Li}_{r}(e^{-w})=\sum_{n=1}^{\infty}\frac{e^{-nw}}{n^{r}}.$
As this expression represents a harmonic sum, we apply Mellin transform and obtain
$\Lambda^{*}(s)=\int_{0}^{\infty}\sum_{n=1}^{\infty}\frac{e^{-nw}}{n^{r}}w^{s-1}dw=\sum_{n=1}^{\infty}\frac{1}{n^{s+r}}\int_{0}^{\infty}e^{-v}v^{s-1}dv=\zeta(s+r)\Gamma(s),$
for $\mathrm{Re}(s)>\max(0,1-r)$. The Gamma function has simple poles at the negative integers, while $\zeta(s+r)$ has a simple pole at s=1rs=1-r. Therefore, by applying Theorem 2.3, we obtain
$\mathrm{Li}_{r}(z)=\Gamma(1-r)w^{r-1}+\sum_{j=0}^{\infty}\frac{(-1)^{j}}{j!}\zeta(r-j)w^{j},$
where
$w=\sum_{\ell=1}^{\infty}\frac{(1-z)^{\ell}}{\ell}.$
By applying these estimates in the converse mapping, we derive an asymptotic expression for $\mathcal{Q}_{f,n}$, given by
$\mathcal{Q}_{f,n}(u)=\frac{e^{n\tau+F(\tau,u)}}{\sqrt{2\pi B}}(1+O(\tau^{2r/7-1})).$ (2.13)
Recall the moment-generating function
$M_{n}(\theta)=\mathbb{E}(e^{(\varpi_{f,n}-\mu_{f,n})\theta/\nu_{f,n}})=\exp\left(-\frac{\mu_{f,n}\theta}{\nu_{f,n}}\right)\mathbb{E}(e^{\theta\varpi_{f,n}/\nu_{f,n}})=\exp\left(-\frac{\mu_{f,n}\theta}{\nu_{f,n}}\right)\frac{\mathcal{Q}_{n}(e^{1/\nu_{f,n}})}{\mathcal{Q}_{n}(1)}.$
Next, by applying the Taylor expansion of $M_{n}(\theta)$, we obtain
$M_{n}(\theta)=\exp\left(\frac{\theta^{2}}{2}+O(n^{-2r/7(r+2)+1/(r+2)})\right),$
thereby establishing the central limit theorem for $\varpi_{f,n}$.
3. The Dirichlet series
In this section, we examine the asymptotic behaviour of the Dirichlet series twisted by the shifted divisor coefficient, which plays a crucial role in deriving the main term. By substituting the expression for $\Delta_{f}(n)$ into the Dirichlet series given in (2.10), we arrive at
$\mathcal{D}_{f}(\Delta;s)\coloneqq\sum_{n=1}^{\infty}\frac{f(n+1)-f(n)}{n^{s}}\eqqcolon A_{f}(s)-D_{f}(s),$ (3.1)
where
$A_{f}(s)=\sum_{n=1}^{\infty}\frac{f(n+1)}{n^{s}}$ (3.2)
is the Dirichlet series with a shifted coefficient and
$D_{f}(s)=\sum_{n=1}^{\infty}\frac{f(n)}{n^{s}},$ (3.3)
for $\mathrm{Re}(s)>\alpha$ and $\alpha\in\mathbb{R}_{>0}$.
Note that for a class of multiplicative functions $f:\mathbb{N}\to\mathbb{N}$ with period 1, i.e. $f(n+1)=f(n)$ for all values of $n$, the analytic properties of $A_{f}(s)$ are closely analogous to those of $D_{f}(s)$. One notable example in this context is the Dirichlet character $\chi(n)$ with period 1. However, determining the analytic behaviour of $A_{f}(s)$ (such as its poles and residues) directly from the Dirichlet series $D_{f}(s)$ for general multiplicative functions $f$ remains a difficult problem.
The study of the distribution and growth of the gap function $(|f(n+1)-f(n)|)_{n}$ presents an intriguing area of research, leading to significant insights into the distribution of multiplicative functions at primes. In this context, notable contributions have been made by O. Klurman [18] and A. Mangerel [24]. Using the properties of correlations of “pretentious” functions, Klurman [18] counts the number of ways to write $n=a+b$, where $a,b$ belongs to some multiplicative subsets of $\mathbb{N}$. Recently, Mangerel [24] has explored this type of result in further detail. However, the main result of this paper takes a different approach, connecting instead with counting problems of the type described in (1.4).
In particular, the Ramanujan sum $c_{m}(n)$ possesses arithmetic identities with a broad class of multiplicative functions. Furthermore, it can be interpreted as a character, providing significant leverage to deduce the properties of $A_{f}(s)$ from $D_{f}(s)$ for those multiplicative functions $f:\mathbb{N}\to\mathbb{N}$ that exhibit an identity with the Ramanujan sum.
For the case of generalised divisor function $f(n)=\sigma_{r}(n)$, we use the identity (see for instance [19]) for fixed $r\geq 1$,
$\sigma_{r}(n)=\zeta(r+1)n^{r}\sum_{m=1}^{\infty}\frac{c_{m}(n)}{m^{r+1}},$ (3.4)
with the Ramanujan sum given in (2.1).
Note that the case $0\leq r<1$ poses different challenges within the framework of our analysis. Observe that when $f(n)=n^{\beta}$ for $\beta\in\mathbb{R}$, it simplifies to the result in [22, Theorem 2].
The function $f$, considered in this paper satisfies an identity involving the Ramanujan sum $c_{m}(n)$. These relations are crucial for handling the Dirichlet series $A_{f}(s)$ with shifted a coefficient. Moreover, our proof can be extended to encompass any multiplicative functions that exhibit similar identities with the Ramanujan sum. Notably, this method also applies to the von Mangoldt function $\Lambda(n)$. Although $\Lambda(n)$ is not a multiplicative function, it satisfies a similar identity involving the Ramanujan sum. Consequently, the arguments presented in this article are sufficient to establish a central limit theorem, analogous to Theorem 2.2, for prime-detecting function that has an identity with $c_{m}(n)$.
Remark 3.1.
In this context, it is important to observe that the Ramanujan sum $c_{m}(n)$ is a sign-changing function, which implies that integer partitions involving the Ramanujan sum cannot be expressed in terms of (1.1). Nevertheless, an analogue of Theorem 2.2 can be established for the Ramanujan sum, subject with certain restrictions on $m$.
For Dirichlet character $\chi$ modulo $m$, the generalised Gauss sum is defined as
$c^{\prime}_{\chi}(n)\coloneqq\sum_{\begin{subarray}{c}b(\bmod{m})\\ (b,m)=1\end{subarray}}\chi(b)e\left(\frac{bn}{m}\right).$ (3.5)
Before analysing the Dirichlet series $A_{f}(s)$, we establish some prerequisites.
Lemma 3.2.
Let $c_{m}(n)$ denote the Ramanujan sum for $m\geq 1$ and $n\geq 1$. Then the following identity holds
$c_{m}(n+1)=\frac{\mu(m)}{\varphi(m)}c_{m}(n)+\frac{1}{\varphi(m)}\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\tau(\chi)c^{\prime}_{\bar{\chi}}(n),$
where $\tau(\chi)=c^{\prime}_{\chi}(1)$ has defined in (3.5).
Proof.
Considering the Fourier expansion of the additive character (see, for instance, [17, §3.4]), we write
$e\left(\frac{b}{m}\right)=\frac{1}{\varphi(m)}\sum_{\begin{subarray}{c}\chi(\bmod{m})\end{subarray}}\bar{\chi}(b)\tau(\chi),\quad\text{if $(b,m)=1$},$ (3.6)
where $\tau(\chi)=\sum_{b(\bmod{m})}\chi(b)e(b/m)$ denotes the Gauss sum associated with the Dirichlet character $\chi$. We simplify the shifted Ramanujan sum as follows
$c_{m}(n+1)=\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}e\left(\frac{b(n+1)}{m}\right)=$ $\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}e\left(\frac{bn}{m}\right)e\left(\frac{b}{m}\right)$
$=$ $\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}e\left(\frac{bn}{m}\right)\left(\frac{1}{\varphi(m)}\sum_{\chi(\bmod{m})}\bar{\chi}(b)\tau(\chi)\right)$
$=$ $\frac{1}{\varphi(m)}\sum_{\chi(\bmod{m})}\tau(\chi)\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}\bar{\chi}(b)e\left(\frac{bn}{m}\right),$ (3.7)
where we substituted (3.6) into the second line. We shall isolate the terms associated with principal characters as these are considerably simpler. Consequently, from the preceding expression, we derive
$c_{m}(n+1)=$ $\frac{1}{\varphi(m)}\left(\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi=\chi_{0}\end{subarray}}\tau(\chi)\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}\bar{\chi}(b)e\left(\frac{bn}{m}\right)+\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\tau(\chi)\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}\bar{\chi}(b)e\left(\frac{bn}{m}\right)\right)$
$=$ $\frac{\mu(m)}{\varphi(m)}c_{m}(n)+\frac{1}{\varphi(m)}\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\tau(\chi)c^{\prime}_{\bar{\chi}}(n),$
thereby concluding the lemma. ∎
In the previous lemma, we transformed the shifted Ramanujan sum into a multiplicative structure by using the character sum. This approach facilitates the analysis of the Dirichlet series that arises in the preceding argument.
Lemma 3.3.
For a fixed integer $r\geq 1$, consider the double Dirichlet series
$D_{1}(s,r)\coloneqq$ $\sum_{m=1}^{\infty}\frac{\mu(m)}{\varphi(m)m^{r+1}}\sum_{n=1}^{\infty}\frac{c_{m}(n)}{n^{s}},$ (3.8)
where $c_{m}(n)$ denotes the Ramanujan sum. The following identity then holds for $\mathrm{Re}(s)>1$:
$D_{1}(s,r)=\frac{\zeta(s)}{\zeta(s+r+1)}K_{r}(s),$
where
$K_{r}(s)=C(r)\prod_{p}\left(1+A_{p,r}(s)\right)$ (3.9)
with
$A_{p,r}(s)=-\frac{\frac{1}{p^{r+1}(p-1)}}{1+\frac{1}{p^{r+1}(p-1)}}\left(\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\frac{1}{p^{2s+2r+2}}+\cdots-\frac{1}{p^{(k+1)(s+r+1)}}+\cdots\right)$
and
$C(r)\coloneqq\prod_{p}\left(1+\frac{1}{p^{r+1}(p-1)}\right).$ (3.10)
Proof.
Using relation (3.3) of [17, §3.2], the following expression simplifies as
$\frac{c_{m}(n)}{\varphi(m)}=\frac{1}{\varphi(m)}\frac{\varphi(m)\mu\left({m}/{(m,n)}\right)}{\varphi\left({m}/{(m,n)}\right)}=\frac{\mu\left({m}/{(m,n)}\right)}{\varphi\left({m}/{(m,n)}\right)}.$
By rearranging the terms in (3.8) and substituting the above relation, we obtain
$D_{1}(s,r)=\sum_{m=1}^{\infty}\frac{\mu(m)}{m^{r+1}}\sum_{n=1}^{\infty}\frac{\mu({m}/{(m,n)})}{\varphi\left({m}/{(m,n)}\right)n^{s}}.$ (3.11)
Let $\delta=(m,n)$, then we may write $m=k\delta$ and $n=\ell\delta$. Consequently, upon making the change of variables in (3.11), we obtain
$D_{1}(s,r)=$ $\sum_{\delta\geq 1}\mathop{\sum\sum}_{(m,n)=\delta}\frac{\mu(m)}{m^{r+1}}\frac{1}{n^{s}}\frac{\mu({m}/{(m,n)})}{\varphi\left({m}/{(m,n)}\right)}$
$=$ $\sum_{\delta\geq 1}\mathop{\sum\sum}_{(k,\ell)=1}\frac{\mu(k\delta)}{(k\delta)^{r+1}}\frac{1}{(\ell\delta)^{s}}\frac{\mu(k)}{\varphi(k)}.$
Rearranging the order of summation yields
$D_{1}(s,r)=\sum_{k=1}^{\infty}\frac{\mu^{2}(k)}{\varphi(k)k^{r+1}}\mathop{\sum\sum}_{(\delta,k)=1}\frac{\mu(\delta)}{\delta^{s+r+1}}\sum_{(\ell,k)=1}\frac{1}{\ell^{s}}.$
The sum over $\ell$ equals to
$\zeta(s)\prod_{p|k}\left(1-\frac{1}{p^{s}}\right),$
and the sum over $\delta$ equals to
$\frac{1}{\zeta(s+r+1)}\prod_{p|k}\left(1-\frac{1}{p^{s+r+1}}\right)^{-1}.$
By rearranging the sum over $k$ and invoking multiplicativity, we deduce that
$D_{1}(s,r)=\frac{\zeta(s)}{\zeta(s+r+1)}\prod_{p}\left(1+\frac{1}{p^{r+1}(p-1)}\frac{1-p^{-{s}}}{1-p^{-(s+r+1)}}\right)=\frac{\zeta(s)}{\zeta(s+r+1)}K_{r}(s).$ (3.12)
Observe that as $\mathrm{Re}(s)\to\infty$, the factors $1-p^{-s}$ and $1-p^{-(s+r+1)}$ tend to 1 for any fixed integer $r\geq 1$. In contrast, the factor ${1}/{p^{r+1}(p-1)}$ remains unchanged, as it is independent of $s$. Hence, we normalise the Euler product $K_{r}(s)$ as
$K_{r}(s)=$ $\prod_{p}\frac{\left(1+\frac{1}{p^{r+1}(p-1)}\frac{1-p^{-{s}}}{1-p^{-(s+r+1)}}\right)}{1+\frac{1}{p^{r+1}(p-1)}}\left(1+\frac{1}{p^{r+1}(p-1)}\right)$
$=$ ${C(r)}\prod_{p}\frac{\left(1+\frac{1}{p^{r+1}(p-1)}(1+B_{p,r}(s))\right)}{1+\frac{1}{p^{r+1}(p-1)}},$
where $C(r)$ is given in (3.10), and
$B_{p,r}(s)=\frac{1-\frac{1}{p^{s}}}{1-\frac{1}{p^{s+r+1}}}-1=$ $\left(1-\frac{1}{p^{s}}\right)\left(1+\frac{1}{p^{s+r+1}}+\frac{1}{p^{2(s+r+1)}}+\cdots\right)-1$
$=$ $-\left(\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\cdots-\frac{1}{p^{(k+1)(s+r+1)}}+\cdots\right).$ (3.13)
In this context, it is worth noting that the decomposition argument employed herein was adopted from [1, §3]. Thus, the Euler product in (3.12) can be expressed as
$K_{r}(s)=C(r)\prod_{p}(1+A_{p,r}(s)),$
where
$A_{p,r}(s)=$ $\frac{1+\left(\frac{1}{p^{r+1}(p-1)}\left(1+B_{p,r}(s)\right)\right)}{1+\frac{1}{p^{r+1}(p-1)}}-1$
$=$ $-\frac{\frac{1}{p^{r+1}(p-1)}}{1+\frac{1}{p^{r+1}(p-1)}}\left(\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\frac{1}{p^{2s+2r+2}}+\cdots-\frac{1}{p^{(k+1)s+r+1}}+\cdots\right).$
By (3), the above assertion follows, thereby concluding the lemma. ∎
Remark 3.4.
Note that the constant $C(r)$ defined in (3.10) may be regarded as the generalised totient summatory constant. In particular, when $r=1$, we obtain
$C(1)=\prod_{p}\left(1+\frac{1}{p^{2}(p-1)}\right)=1.339784\ldots,$
a result that can be deduced from the well-known Landau totient constant
$A=\zeta(2)C(1)=\zeta(2)\prod_{p}\left(1+\frac{1}{p^{2}(p-1)}\right)=2.20386\ldots.$
Let $s=\sigma+it$ with $\sigma,t\in\mathbb{R}$. We have
$\left|1-\frac{1}{p^{s+r+1}}\right|\geq 1-\frac{1}{2^{\sigma+r+1}},\quad\mathrm{and}\quad\left|\frac{1}{p^{s}}\right|=\frac{1}{p^{\sigma}},$
then
$|B_{p,r}|=$ $\left|\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\frac{1}{p^{2s+2r+2}}+\cdots-\frac{1}{p^{(k+1)s+r+1}}+\cdots\right|<$ $\frac{\frac{1}{p^{\sigma}}}{1-\frac{1}{p^{\sigma+r+1}}}.$
Additionally, for a fixed integer $r\geq 1$,
${\frac{1}{p^{r+1}(p-1)}}\left({1+\frac{1}{p^{r+1}(p-1)}}\right)^{-1}<\frac{1}{p^{r+1}}.$
so that
$\sum_{p}|A_{p,r}(s)|<\frac{2^{\sigma+r+1}}{2^{\sigma+r+1}-1}\sum_{p}\frac{1}{p^{\sigma+r+1}}<\infty.$
It follows that the Euler product $K_{r}(s)$ converges absolutely (see, for instance, [31, §1.42]) and defines an analytic function on the half-plane $\mathrm{Re}(s)>1$. Furthermore, for any fixed $\sigma_{0}>-r$,
$|K_{r}(s)|\leq C(r)\prod_{p}\left(1+|A_{p,r}(s)|\right)\leq C(r)\prod_{p}\left(1+\frac{2^{\sigma_{0}+r+1}}{(2^{\sigma_{0}+r+1}-1)p^{\sigma_{0}+r+1}}\right).$
Hence, $K_{r}(s)$ is uniformly bounded on the half-plane $\mathrm{Re}(s)>\sigma_{0}$. Since it is analytic for $\mathrm{Re}(s)>1$, the Dirichlet series $D_{1}(s,r)$ admits a meromorphic continuation in this region. Moreover, because $\zeta(s)$, it has a simple pole at point $s=1$ and $1/\zeta(s+r+1)$ has no zeros on the half-plane where $\mathrm{Re}(s)+r>0$, it follows that $D_{1}(s,r)$ extends analytically to the region $\mathrm{Re}(s)+r>0$ and $\mathrm{Re}(s)>1$ except for a simple pole at $s=1$. Assuming the Riemann Hypothesis, the function may be analytically continued to the half-plane $\mathrm{Re}(s)>-r-\frac{1}{2}$.
From Lemma 3.3 and the foregoing discussion, we conclude the following result.
Proposition 3.5.
For a fixed integer $r\geq 1$ and $s=\sigma+it$ with $\sigma,t\in\mathbb{R}$, the Dirichlet series $D_{1}(s,r)$ has analytic continuation to the half plane for $\mathrm{Re}(s)>-r$ and $\mathrm{Re}(s)>1$, with an exception of a pole at the point $s=1$. Furthermore, this function has analytic continuation to the half-plane $\mathrm{Re}(s)>-r-\frac{1}{2}$ if and only if the Riemann hypothesis is true.
Observe that
$c^{\prime}_{\chi}(n)=\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}\chi(b)e\left(\frac{bn}{m}\right)=\sum_{\begin{subarray}{c}b\bmod{m}\end{subarray}}\chi_{0}(b)\chi(b)e\left(\frac{bn}{m}\right)=\sum_{\begin{subarray}{c}b\bmod{m}\end{subarray}}\chi_{1}(b)e\left(\frac{bn}{m}\right)=c^{\prime}_{\chi_{1}}(n),$ (3.14)
where $\chi_{1}$ is a Dirichlet character modulo $m$ defined by $\chi_{1}=\chi_{0}\chi$. Note that in Lemma 3.2, the Dirichlet character $\chi$ is not necessarily a primitive character. Therefore, we further induce it by $\chi^{*}$ modulo $m^{*}$, where $\chi^{*}$ is a primitive Dirichlet character such that $m^{*}|m$.
11footnotetext: Note that $\chi^{*}_{0}\chi_{0}$ is a principal character mod $m$ because $m^{*}|m$.
Then we can write $\chi=\chi^{*}_{0}\chi^{*}$ for principal character $\chi^{*}_{0}$. Then the generalised Gauss sum can be written as
$c_{{\chi}}(n)=$ $\sum_{\begin{subarray}{c}b\bmod{m}\end{subarray}}{\chi}_{0}(b)\chi^{*}(b)e\left(\frac{bn}{m}\right)=\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}\chi^{*}(b)e\left(\frac{bn}{m}\right)$
$=$ $\frac{\varphi(m)}{\varphi(m^{*})}\sum_{\begin{subarray}{c}b\bmod{m^{*}}\\ (b,m^{*})=1\end{subarray}}\chi^{*}(b)e\left(\frac{bn}{m^{*}}\right)=\frac{\varphi(m)}{\varphi(m^{*})}c^{\prime}_{\chi^{*}}(n),$ (3.15)
where we have used [2, Theorem 5.33(a)] and (3.5) in the last step.
Lemma 3.6.
Let $r>1$ be fixed, and for a Dirichlet character $\chi$ modulo $m$, define
$D_{2}(s,r)\coloneqq\sum_{m=1}^{\infty}\frac{1}{\varphi(m)m^{r+1}}\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\tau(\chi)\sum_{n=1}^{\infty}\frac{c^{\prime}_{\bar{\chi}}(n)}{n^{s}}.$ (3.16)
Let $\mathrm{Re}(s)=\sigma$, then the following inequality
$|D_{2}(s,r)|\leq\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)}{\zeta(2(\sigma+r))}\zeta(r)E_{r}(\sigma)$
holds for $\sigma>1$, where
$E_{r}(\sigma)=C^{\prime}(r)\prod_{p}(1+J_{p,r}(\sigma))$ (3.17)
with
$J_{p,r}(\sigma)=-\frac{\frac{1-p^{-r}}{p^{r+1}}\left(\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)}{1+\frac{1-p^{-r}}{p^{r+1}}}$
and
$C^{\prime}(r)\coloneqq\prod_{p}\left(1+\frac{1}{p^{r+1}}\left(1-\frac{1}{p^{r}}\right)\right).$ (3.18)
Proof.
Let $\chi$ be a Dirichlet character modulo $m$, induced by a primitive character $\chi^{*}$ modulo $m^{*}$, where $m^{*}|m$. From (3.14), it is evident that $c^{\prime}_{\bar{\chi}}(n)=c^{\prime}_{\bar{\chi}_{1}}(n)$ with $\bar{\chi}_{1}=\chi_{0}\bar{\chi}=\chi_{0}\chi^{*}_{0}\bar{\chi}^{*}=\chi_{0}\bar{\chi}^{*}$.
Beginning with the simplification of the coefficient of the Dirichlet series, and applying Theorem 9.10 and Theorem 9.12 of [26], we obtain
$c^{\prime}_{\bar{\chi}}(n)c^{\prime}_{{\chi}}(1)=$ $\chi^{*}\left(\frac{n}{(m,n)}\right)\bar{\chi}^{*}\left(\frac{m/(m,n)}{m^{*}}\right)\mu\left(\frac{m/(m,n)}{m^{*}}\right)\frac{\varphi(m)}{\varphi(m/(m,n))}\mu\left(\frac{m}{m^{*}}\right)\chi^{*}\left(\frac{m}{m^{*}}\right)c^{\prime}_{\bar{\chi}^{*}}(1)c^{\prime}_{{\chi^{*}}}(1)$
$=$ $m^{*}\chi^{*}\left(\frac{n}{(m,n)}\right)\bar{\chi}^{*}\left(\frac{m/(m,n)}{m^{*}}\right)\mu\left(\frac{m/(m,n)}{m^{*}}\right)\frac{\varphi(m)}{\varphi(m/(m,n))}\mu\left(\frac{m}{m^{*}}\right)\chi^{*}\left(\frac{m}{m^{*}}\right).$
Let $m=\delta m^{*}k$ and $n=\delta\ell$ with $(m,n)=\delta$, so that $(m^{*}k,\ell)=1$. With this change of variables, the above expression becomes
$\frac{c^{\prime}_{\bar{\chi}}(n)c^{\prime}_{{\chi}}(1)}{\varphi(m)}=$ $m^{*}\chi^{*}\left(\ell\right)\bar{\chi}^{*}\left(k\right)\mu\left(k\right)\frac{1}{\varphi(m^{*}k)}\mu\left(\delta k\right)\chi^{*}\left(\delta k\right)$
$=$ $m^{*}\chi^{*}\left(\ell\right)\bar{\chi}^{*}\left(k\right)\mu\left(k\right)\frac{1}{\varphi(m^{*}k)}\mu\left(\delta\right)\mu\left(k\right)\chi^{*}\left(\delta\right)\chi^{*}\left(k\right)$
$=$ $m^{*}\chi^{*}\left(\ell\right)\mu^{2}\left(k\right)\frac{1}{\varphi(m^{*}k)}\mu\left(\delta\right)\chi^{*}\left(\delta\right).$ (3.19)
Rearranging the terms in (3.16), and substituting (3) with the change of variable, we obtain
$D_{2}(s,r)=$ $\sum_{m=1}^{\infty}\frac{1}{m^{r+1}}\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\sum_{n=1}^{\infty}\frac{c^{\prime}_{\chi}(1)c^{\prime}_{\bar{\chi}}(n)}{\varphi(m)n^{s}}$
$=$ $\sum_{\delta\geq 1}\mathop{\sum\sum}_{\begin{subarray}{c}k,\ell\geq 1\\ (\ell,k)=1\end{subarray}}\sum_{m^{*}\geq 1}\frac{\varphi(m^{*}k\delta)}{\varphi(m^{*})\varphi(m^{*}k)}\sum_{\chi^{*}(\bmod{m^{*}})}\frac{m^{*}\mu(\delta k)}{(m^{*}\delta k)^{r+1}}\chi^{*}(\delta k)\frac{\mu(k)\chi^{*}(\ell)\bar{\chi}^{*}(k)}{(\ell\delta)^{s}}.$
Furthermore, for $(m^{*},k)=1$ and $(\delta,k)=1$, we have
$\frac{\varphi(m)}{\varphi(m^{*}k)\varphi(m^{*})}=\frac{\varphi(m^{*}k\delta)}{\varphi(m^{*}k)\varphi(m^{*})}=\frac{\varphi(\delta)}{\varphi(m^{*})}.$
Rearranging the order of summation, we obtain
$D_{2}(s,r)=\sum_{k\geq 1}\frac{\mu^{2}(k)}{k^{r+1}}\sum_{\begin{subarray}{c}m^{*}\geq 1\\ ({m^{*}},k)=1\end{subarray}}\frac{1}{\varphi(m^{*}){m^{*}}^{r}}\sum_{\begin{subarray}{c}\chi^{*}(\bmod m^{*})\end{subarray}}\sum_{\begin{subarray}{c}\delta\geq 1\\ (\delta,k)=1\end{subarray}}\frac{\mu(\delta)\chi^{*}(\delta)\varphi(\delta)}{\delta^{s+r+1}}\sum_{\begin{subarray}{c}\ell\geq 1\\ (\ell,k)=1\end{subarray}}\frac{\chi^{*}(\ell)}{\ell^{s}}.$
The sum over $\ell$ equals to
$L(s,\chi^{*})\prod_{\begin{subarray}{c}p\mid k\\ p\nmid m^{*}\end{subarray}}\left(1-\frac{\chi^{*}(p)}{p^{s}}\right),$ (3.20)
and the sum over $\delta$ equals to
$\prod_{\begin{subarray}{c}p\nmid k\\ p\nmid m^{*}\end{subarray}}\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right).$ (3.21)
Observe that, by utilising a decomposition argument analogous to that employed in the preceding lemma, the expression above may be equivalently simplified to
$\frac{L(s+r+1,\chi^{*})}{L(s+r,\chi^{*})}\prod_{\begin{subarray}{c}p|k\\ p\nmid m^{*}\end{subarray}}(1+B_{p,r}(s,\chi^{*}))\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right)^{-1},$
where
$B_{p,r}(s,\chi^{*})=-\frac{\left(\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r)+1}}-\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r+1)}}+\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r)+1}}-\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r+1)}}+\cdots\right)}{\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}-\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r)+1}}+\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r+1)}}-\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r)+1}}+\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r+1)}}-\cdots\right)}.$ (3.22)
Note that the stipulation $p\nmid m^{*}$ in the Euler products mentioned above is superfluous, as $\chi^{*}(p)=0$ for every prime $p|m^{*}$. Set
$A_{r}(s,\chi^{*})=\prod_{p|k}(1+B_{p,r}(s,\chi^{*}))\left(1-\frac{\chi^{*}(p)}{p^{s}}\right)\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right)^{-1}.$ (3.23)
Thus, the average over the characters is given by
$K_{r}(s,\chi^{*})\coloneqq\sum_{\chi^{*}(\bmod{m^{*}})}\frac{L(s,\chi^{*})L(s+r+1,\chi^{*})}{L(s+r,\chi^{*})}A_{r}(s,\chi^{*}).$ (3.24)
We begin by estimating the term $A_{r}(s,\chi^{*})$ and establishing its region of convergence. Letting $s=\sigma+it$ with $\sigma,t\in\mathbb{R}$, and recalling (3.22), we have
$|B_{p,r}(s,\chi^{*})|\leq\frac{1}{p^{2\sigma+2r+1}}\frac{1-\frac{1}{2^{\sigma+r+1}}}{1-\frac{1}{2^{\sigma+r}}}<\frac{1}{2p^{2\sigma+2r+1}},$
and consequently,
$\sum_{p|k}|B_{p,r}(s,\chi^{*})|<\frac{1}{2}\sum_{p|k}\frac{1}{p^{2\sigma+2r+1}}<\infty.$
Next, consider the factors
$\left|1-\frac{\chi^{*}(p)}{p^{s}}\right|\leq\left|1+\frac{1}{p^{s}}\right|=1+\frac{1}{p^{\sigma}}\quad\text{and}\quad$ $\left|1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right|\geq 1-\frac{1}{2^{\sigma+r+1}}.$
It follows that
$|A_{r}(s,\chi^{*})|\leq$ $\prod_{p|k}|1+B_{p,r}(s,\chi^{*})|\left|1+\frac{1}{p^{s}}\right|\left|1+\frac{1}{p^{s+r+1}}-\frac{1}{p^{s+r}}\right|^{-1}$
$\leq$ $\prod_{p|k}\left(1+\frac{1}{2p^{2\sigma+2r+1}}\right)\left(1+\frac{1}{p^{\sigma}}\right)\left(1-\frac{1}{2^{(\sigma+r+1)}}\right)^{-1}$
$\leq$ $\left(1-\frac{1}{2^{(\sigma+r+1)}}\right)^{-1}\prod_{p|k}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}+\frac{1}{2p^{3\sigma+2r+1}}\right).$ (3.25)
Therefore, $|A_{r}(s,\chi^{*})|$ is uniformly bounded in the half-plane $\mathrm{Re}(s)>1$. Moreover, since
$\frac{\zeta(2\sigma)}{\zeta(\sigma)}\leq|L(s,\chi^{*})|\leq\zeta(\sigma)$
for $\mathrm{Re}(s)=\sigma>1$ (see for instance [5, Lemma 2.1]), we may bound the sum in (3.24) by
$|K_{r}(s,\chi^{*})|\leq$ $\left|\sum_{\chi^{*}(\bmod{m^{*}})}\frac{L(s,\chi^{*})L(s+r+1,\chi^{*})}{L(s+r,\chi^{*})}A_{r}(s,\chi^{*})\right|$
$\leq$ $\varphi(m^{*})\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)}{\zeta(2(\sigma+r))}\prod_{p|k}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\left(1-\frac{1}{2^{\sigma+r+1}}\right)^{-1}.$
Ignoring the higher-order terms and inserting the expression above, we obtain
$|D_{2}(s,r)|\leq\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)}{\zeta(2(\sigma+r))}\sum_{k\geq 1}\frac{\mu^{2}(k)}{k^{r+1}}\prod_{p|k}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\sum_{\begin{subarray}{c}m^{*}\geq 1\\ (m^{*},k)=1\end{subarray}}\frac{1}{{m^{*}}^{r}}.$
The sum over $m^{*}$ equals to
$\zeta(r)\prod_{p|k}\left(1-\frac{1}{p^{r}}\right).$
Hence, we deduce that
$|D_{2}(s,r)|\leq\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)\zeta(r)}{\zeta(2(\sigma+r))}\prod_{p}\left(1+\frac{1-p^{-r}}{p^{r+1}}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\right),$
and set
$E_{r}(\sigma)=\prod_{p}\left(1+\frac{1-p^{-r}}{p^{r+1}}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\right).$
Furthermore, the Euler product can be expressed in the form
$E_{r}(\sigma)=C^{\prime}(r)\prod_{p}\frac{\left(1+\frac{1-p^{-r}}{p^{r+1}}(1+E_{p,r}(\sigma))\right)}{1+\frac{1-p^{-r}}{p^{r+1}}},$
where $C^{\prime}(r)$ is defined in (3.18) and
$E_{p,r}(\sigma)=-\left(\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right).$
From our analysis of $B_{p,r}(s,\chi^{*})$, it follows that $E_{p,r}(\sigma)$ converges absolutely for $\sigma>1$, thereby completing the proof of the lemma. ∎
3.1. Special values of $D_{2}(s,r)$
It is worth noting that the work of Ramaré [27] can be used to handle the Euler products over arithmetic progressions in (3.20) and (3.21) and for deriving an exact formula. However, the upper bound established for this Dirichlet series in the preceding lemma suffices for the purposes of this paper. Here, we determine the asymptotic behaviour of $D_{2}(s,r)$ in the case of a particular Dirichlet character $\chi_{4}$.
Define
$\chi_{4}(n)\coloneqq\begin{cases}1&\text{if $n\equiv 1(\bmod{4})$}\\ -1&\text{if $n\equiv 3(\bmod{4})$}\\ 0&\text{if $2|n$}.\end{cases}$ (3.26)
Note that the Dirichlet $L$-function $L(s,\chi_{4})=\beta(s)$, the Dirichlet $\beta$-function, is given by
$L(s,\chi_{4})=\sum_{n=1}^{\infty}\frac{\chi_{4}(n)}{n^{s}}=\sum_{\ell=0}^{\infty}\frac{(-1)^{\ell}}{(2\ell+1)^{s}}.$
Since $m^{*}=4$ and $(m^{*},k)=1$, it is evident from the proof of Lemma 3.6 that
$D_{2}(s,r;\chi_{4})=\frac{L(s,\chi_{4})L(s+r+1,\chi_{4})}{L(s+r,\chi_{4})}\sum_{k=1}^{\infty}\frac{\mu^{2}(k)}{k^{r+1}}A_{r}(s,\chi_{4})$
where
$A_{r}(s,\chi_{4})=$ $\prod_{\begin{subarray}{c}p|k\end{subarray}}(1-B_{p,r}(s,\chi_{4}))\left(1+\frac{\chi_{4}(p)}{p^{s+r+1}}-\frac{\chi_{4}(p)}{p^{s+r}}\right)^{-1}$
$=$ $\prod_{\begin{subarray}{c}p|k\\ p\equiv 1(\bmod{4})\end{subarray}}(1-B_{p,r}(s))\left(1-\frac{1}{p^{s}}\right)\left(1+\frac{1}{p^{s+r+1}}-\frac{1}{p^{s+r}}\right)^{-1}$
$\quad\quad\quad\quad\times\prod_{\begin{subarray}{c}p|k\\ p\equiv 3(\bmod{4})\end{subarray}}(1-B_{p,r}(s))\left(1+\frac{1}{p^{s}}\right)\left(1-\frac{1}{p^{s+r+1}}+\frac{1}{p^{s+r}}\right)^{-1},$
where $B_{p,r}(s)$ is given in (3.22). Thus, $D_{2}(s,r;\chi_{4})$ equals to
$\frac{L(s,\chi_{4})L(s+r+1,\chi_{4})}{L(s+r,\chi_{4})}\prod_{p\equiv 1(\bmod{4})}\left(1+\frac{(1+p^{-s})}{p^{r+1}}\left(1+\frac{1}{p^{s+r+1}}-\frac{1}{p^{s+r}}\right)^{-1}(1-B_{p,r}(s))\right)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\prod_{p\equiv 3(\bmod{4})}\left(1+\frac{(1-p^{-s})}{p^{r+1}}\left(1-\frac{1}{p^{s+r+1}}+\frac{1}{p^{s+r}}\right)^{-1}(1-B_{p,r}(s))\right)$
$=$ $\frac{\beta(s)\beta(s+r+1)}{\beta(s+r)}\prod_{p\equiv 1(\bmod{4})}\left(1+\frac{(1+p^{-s})}{p^{r+1}}\left(1+\frac{1}{p^{s+r+1}}-\frac{1}{p^{s+r}}\right)^{-1}(1-B_{p,r}(s))\right)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\prod_{p\equiv 3(\bmod{4})}\left(1+\frac{(1-p^{-s})}{p^{r+1}}\left(1-\frac{1}{p^{s+r+1}}+\frac{1}{p^{s+r}}\right)^{-1}(1-B_{p,r}(s))\right),$
where $\beta(s)$ denotes the Dirichlet $\beta$ function and we can obtain a decomposed product as before. Note that for $\mathrm{Re}(s)=1$ the above Euler product over arithmetic progression reduced to specific constants.
Lemma 3.7.
For a fixed integer $r>1$, the Dirichlet series
$A_{\sigma_{r}}(s)=\sum_{n=1}^{\infty}\frac{\sigma_{r}(n+1)}{n^{s}}=$ $\zeta(r+1)\sum_{i=0}^{r}\binom{r}{i}D_{1}(s-i,r)+\zeta(r+1)\sum_{i=0}^{r}\binom{r}{i}D_{2}(s-i,r),$
converges for $\mathrm{Re}(s)>1$, where $D_{1}(s,r)$ and $D_{2}(s,r)$ are defined in (3.8) and (3.16), respectively.
Proof.
Employing the identity (3.4), we express
$A_{\sigma_{r}}(s)=\sum_{n=1}^{\infty}\frac{\sigma_{r}(n+1)}{n^{s}}=\zeta(r+1)\sum_{n=1}^{\infty}\frac{(n+1)^{r}}{n^{s}}\sum_{m=1}^{\infty}\frac{c_{m}(n+1)}{m^{r+1}}.$
Given that $r\geq 1$ is fixed, and under the conditions $\mathrm{Re}(s)>1$, and $\mathrm{Re}(s)+r>1$, one may interchange the order of summation, thereby obtaining
$\zeta(r+1)\sum_{m=1}^{\infty}\frac{1}{m^{r+1}}\sum_{n=1}^{\infty}\frac{c_{m}(n+1)}{n^{s}}(n+1)^{r}=\zeta(r+1)\sum_{i=0}^{r}\binom{r}{i}\sum_{m=1}^{\infty}\frac{1}{m^{r+1}}\sum_{n=1}^{\infty}\frac{c_{m}(n+1)}{n^{s-i}}.$ (3.27)
Now applying Lemma 3.2 yields,
$\sum_{m=1}^{\infty}\frac{1}{m^{r+1}}\sum_{n=1}^{\infty}\frac{c_{m}(n+1)}{n^{s-i}}=$ $\sum_{m=1}^{\infty}\frac{1}{\varphi(m)m^{r+1}}\sum_{n=1}^{\infty}\frac{1}{n^{s-i}}\left(\mu(m)c_{m}(n)+\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}c_{\chi}(1)c^{\prime}_{\bar{\chi}}(n)\right)$
$=$ $\sum_{m=1}^{\infty}\frac{\mu(m)}{\varphi(m)m^{r+1}}\sum_{n=1}^{\infty}\frac{c_{m}(n)}{n^{s-i}}+\sum_{m=1}^{\infty}\frac{1}{\varphi(m)m^{r+1}}\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\tau(\chi)\sum_{n=1}^{\infty}\frac{c^{\prime}_{\bar{\chi}}(n)}{n^{s-i}}$
$=$ $D_{1}(s-i,r)+D_{2}(s-i,r).$
By applying Lemma 3.3 and Lemma 3.6, in conjunction with (3.27), we thereby complete the proof. ∎
We conclude this section by recalling that (see, for instance, [31, §1.3]), for $\mathrm{Re}(s)>1$ or $\mathrm{Re}(s)>r+1$,
$D_{\sigma_{r}}(s)=\sum_{n=1}^{\infty}\frac{\sigma_{r}(s)}{n^{s}}=\zeta(s)\zeta(s-r),$ (3.28)
for fixed $r$.
4. The Integral Asymptotics
As outlined in Section 2.1, we have two main steps to establish the central limit theorem. First, to employ the saddle-point method, we analyse the asymptotic behaviour of the integrands
$\Omega_{1,f}(n)\coloneqq\frac{e^{n\tau}}{2\pi}\int_{|\theta|\leq\theta_{n}}\exp(in\theta+F(\tau+i\theta,u))d\theta,$ (4.1)
and
$\Omega_{2,f}(n)\coloneqq\frac{e^{n\tau}}{2\pi}\int_{\theta_{n}<|\theta|\leq\pi}\exp(in\theta+F(\tau+i\theta,u))d\theta,$ (4.2)
with a suitable choice of $\tau=\tau(n,u)$ such that the first-order derivative in (2.9) vanishes.
4.1. Estimate of $\Omega_{1,f}$
From the definition of $F(\gamma,u)$, it is evident that deriving an asymptotic expression for the integrand $\Omega_{1,f}$ necessitates a comprehensive understanding of the analytic properties of the Dirichlet series with shifted coefficient $A_{f}(s)$ and $D_{f}(s)$ corresponding to $f(n)=\sigma_{r}(n)$.
Lemma 4.1.
Let $f$ denote the $r$-th divisor function $\sigma_{r}(n)$ for fixed $r\geq 2$, and let $B^{2}=F_{\gamma\gamma}(\tau,u)$. We then have
$\Omega_{1,f}=\frac{e^{n\tau+F(\tau,u)}}{2\pi}\left(\int_{-\theta_{n}}^{\theta_{n}}\exp\left(-\frac{B^{2}}{2}\theta^{2}\right)d\theta\right)\left(1+O\left(\tau^{2r/7-1}\right)\right),$
where
$\int_{-\theta_{n}}^{\theta_{n}}\exp\left(-\frac{B^{2}}{2}\theta^{2}\right)d\theta=\frac{\sqrt{2\pi}}{B}+O\left(\tau^{-1-3r/7}B^{-2}\exp\left(-\frac{\tau^{-r/7-1}}{2}\right)\right)$
uniformly in $u$.
Proof.
Note that
$F(\tau+i\theta,u)=F(\tau,u)+F_{\gamma}(\tau,u)i\theta-\frac{F_{\gamma\gamma}(\tau,u)}{2}\theta^{2}+O\left(\theta^{3}\sup_{0\leq\theta_{0}\leq\theta}|F_{\gamma\gamma\gamma}(\tau+i\theta_{0},u)|\right)$ (4.3)
Utilising the definition of $\Delta_{f}(n)$ (as given in (2.2)), we subsequently deduce that
$F(\gamma,u)=$ $\sum_{n=1}^{\infty}\Delta_{f}(n)\log(1+ue^{-n\gamma})=-\sum_{n=1}^{\infty}\Delta_{f}(n)\sum_{\ell=1}^{\infty}\frac{(-u)^{\ell}}{\ell}e^{-n\ell\gamma}$
$=$ $-\sum_{n=1}^{\infty}(\sigma_{r}(n+1)-\sigma_{r}(n))\sum_{\ell=1}^{\infty}\frac{(-u)^{\ell}}{\ell}e^{-n\ell\gamma}$
$=$ $-\sum_{n=1}^{\infty}\sigma_{r}(n+1)\sum_{\ell=1}^{\infty}\frac{(-u)^{\ell}}{\ell}e^{-n\ell\gamma}+\sum_{n=1}^{\infty}\sigma_{r}(n)\sum_{\ell=1}^{\infty}\frac{(-u)^{\ell}}{\ell}e^{-n\ell\gamma}$
$=$ $-\mathcal{F}_{1}(\gamma,u)+\mathcal{F}_{2}(\gamma,u).$ (4.4)
We begin with the case $\mathcal{F}_{2}(\gamma,u)$ owing to its relative simplicity. For $j\geq 0$, we have
$\frac{\partial^{j}}{\partial\gamma^{j}}\mathcal{F}_{2}(\gamma,u)=(-1)^{j}\sum_{n=1}^{\infty}\sigma_{r}(n)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}e^{-n\gamma\ell}.$ (4.5)
Observe that all the infinite sums above take the form
$\mathfrak{h}_{2,j}(\gamma,u)=\sum_{n=1}^{\infty}\sigma_{r}(n)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}e^{-n\gamma\ell}.$
Let $\mathcal{H}_{2,j}(s,u)$ denote the Mellin transform of $\mathfrak{h}_{2,j}(\gamma,u)$ with respect to $\gamma$. Therefore, by (3.28)
$\mathcal{H}_{2,j}(s,u)=$ $\sum_{n=1}^{\infty}\sigma_{r}(n)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}\int_{0}^{\infty}e^{-n\ell\gamma}\gamma^{s-1}d\gamma$
$=$ $\sum_{n=1}^{\infty}\sigma_{r}(n)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}\Gamma(s)(n\ell)^{-s}$
$=$ $\sum_{n=1}^{\infty}\frac{\sigma_{r}(n)}{n^{s-j}}\sum_{\ell=1}^{\infty}\frac{(-u)^{\ell}}{\ell^{s-j+1}}\Gamma(s)=\zeta(s-j)\zeta(s-j-r)\mathrm{Li}_{s-j+1}(-u)\Gamma(s).$
Note that $\mathcal{H}_{2,j}(s,u)$ converges for $\mathrm{Re}(s)>r+j+2$ and $\mathrm{Re}(s)>j+2$, with its double poles located at the intervals $\frac{3}{2}+j\leq\mathrm{Re}(s)\leq j+2$ and $\frac{3}{2}+r+j\leq\mathrm{Re}(s)\leq r+j+2$ for fixed $r\geq 2$. These poles occur at $s=j+1$ and $s=j+r+1$. Consequently, it is necessary to invoke Theorem 2.3 twice. Thus,
$\mathfrak{h}_{2,j}(\gamma,u)=\zeta(r+1)\mathrm{Li}_{r+2}(-u)\Gamma(j+r+1)\gamma^{-(j+r+1)}$ $+\zeta(1-r)\mathrm{Li}_{2}(-u)\Gamma(j+1)\gamma^{-(j+1)}$
$\quad\quad\quad\quad+O(\gamma^{-(j+3/2)}).$
Thus, inserting the above expression in (4.5), we arrive at
$\frac{\partial^{j}}{\partial\gamma^{j}}\mathcal{F}_{2}(\gamma,u)=(-1)^{j}\mathfrak{h}_{2,j}(\gamma,u)=$ $(-1)^{j}\zeta(r+1)\mathrm{Li}_{r+2}(-u)\Gamma(j+r+1)\gamma^{-(j+r+1)}$
$+(-1)^{j}\zeta(1-r)\mathrm{Li}_{2}(-u)\Gamma(j+1)\gamma^{-(j+1)}+O(\gamma^{-(j+r+3/2)})$
$=$ $(-1)^{j}\zeta(r+1)\mathrm{Li}_{r+2}(-u)\Gamma(j+r+1)\gamma^{-(j+r+1)}+O(\gamma^{-(j+r+3/2)}).$ (4.6)
We now proceed with the computation of $\mathcal{F}_{1}(\gamma,u)$. For $j\geq 0$, we write
$\frac{\partial^{j}}{\partial\gamma^{j}}\mathcal{F}_{1}(\gamma,u)=(-1)^{j}\sum_{n=1}^{\infty}\sigma_{r}(n+1)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}e^{-n\ell\gamma}.$ (4.7)
Similarly, as before, all the infinite sums above take the form
$\mathfrak{h}_{1,j}(\gamma,u)=\sum_{n=1}^{\infty}\sigma_{r}(n+1)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}e^{-n\ell\gamma}.$
Let $\mathcal{H}_{1,j}(s,u)$ denote the Mellin transform of $\mathfrak{h}_{1,j}(\gamma,u)$ with respect to $\gamma$. Thus, we have
$\mathcal{H}_{1,j}(s,u)=$ $\sum_{n=1}^{\infty}\sigma_{r}(n+1)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}\int_{0}^{\infty}e^{-\ell\gamma}\gamma^{s-1}d\gamma$
$=$ $\sum_{n=1}^{\infty}\sigma_{r}(n+1)n^{j}\sum_{\ell=1}^{\infty}(-u)^{\ell}\ell^{j-1}\Gamma(s)(n\ell)^{-s}$
$=$ $\sum_{n=1}^{\infty}\frac{\sigma_{r}(n+1)}{n^{s-j}}\sum_{\ell=1}^{\infty}\frac{(-u)^{\ell}}{\ell^{s-j+1}}\Gamma(s).$
Applying Lemma 3.7, the aforementioned expression becomes
$\mathcal{H}_{1,j}(s,u)=$ $\zeta(r+1)\left(\sum_{i=0}^{r}\binom{r}{i}D_{1}(s-i-j,r)+\sum_{i=0}^{r}\binom{r}{i}D_{2}(s-i-j,r)\right)\mathrm{Li}_{s-j+1}(-u)\Gamma(s).$ (4.8)
Recalling Lemma 3.3, the first term in the above expression yields
$\widetilde{\mathcal{H}}_{1,j}(s,u)=$ $\zeta(r+1)\sum_{i=0}^{r}\binom{r}{i}D_{1}(s-i-j,r)\mathrm{Li}_{s-j+1}(-u)\Gamma(s)$
$=$ $\zeta(r+1)\sum_{i=0}^{r}\binom{r}{i}\frac{\zeta(s-i-j)}{\zeta(s-i-j+r+1)}K_{r}(s-i-j)\mathrm{Li}_{s-j+1}(-u)\Gamma(s),$
where $K_{r}(s)$ is defined in (3.9). Notably, $K_{r}(s)$ is absolutely convergent for $\mathrm{Re}(s)-i-j+r>0$. The only singularity arises from the zeta function at $s=1+i+j$. By applying Theorem 2.3, we obtain
$\widetilde{\mathfrak{h}}_{1,j}(\gamma,u)=$ $\frac{\zeta(r+1)K_{r}(1)}{\zeta(r+2)}\sum_{i=0}^{r}\binom{r}{i}\gamma^{-(i+j+1)}\mathrm{Li}_{i+2}(-u)\Gamma(i+j+1)+O(\gamma^{-r-j-3/2}).$
By Lemma 3.6, we obtain an upper bound for the second term in (4.8) as follows
[…]
$\leq$ $\zeta(r+1)\zeta(r)\sum_{i=0}^{r}\binom{r}{i}\frac{\zeta(\sigma-i-j)\zeta(\sigma-i-j+r)\zeta(\sigma-i-j+r+1)}{\zeta(2(\sigma-i-j+r))}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad E_{r}(\sigma-i-j)\mathrm{Li}_{s-j+1}(-u)\Gamma(s),$
where $\mathrm{Re}(s)=\sigma$ and $E_{r}(\sigma)$ is defined in (3.17). However, triple poles arise from the zeta functions at $\sigma=i+j+1$, $\sigma=i+j+1-r$, and $\sigma=i+j-r$. By invoking Theorem 2.3, we express this as
[…]
$+\zeta(r+1)\zeta(r)\sum_{i=0}^{r}\binom{r}{i}\zeta(1-r)E_{r}(1-r)\mathrm{Li}_{i+2-r}(-u)\Gamma(i+j+1-r)\gamma^{-(i+j+1-r)}$
$+\zeta(r+1)\zeta(r)\sum_{i=0}^{r}\binom{r}{i}{\zeta(-r)E_{r}(-r)}\mathrm{Li}_{i+1-r}(-u)\Gamma(i+j-r)\gamma^{-i-j+r}+O(\gamma^{-r-j-3/2})$
$=$ $\frac{\zeta^{2}(r+1)\zeta(r+2)\zeta(r)}{\zeta(2(r+1))}E_{r}(1)\sum_{i=0}^{r}\binom{r}{i}\mathrm{Li}_{i+2}(-u)\Gamma(i+j+1)\gamma^{-(i+j+1)}+O(\gamma^{-j-r-3/2}).$
11footnotetext: Note that due to the convergence region of $E_{r}(\sigma)$, $E_{r}(-r)$ and $E_{r}(1-r)$ are constants.
[…]
$\leq$ $\frac{\zeta(r+1)K_{r}(1)}{\zeta(r+2)}\sum_{i=0}^{r}\binom{r}{i}\gamma^{-(i+j+1)}\mathrm{Li}_{i+2}(-u)\Gamma(i+j+1)$
$+\frac{\zeta^{2}(r+1)\zeta(r+2)\zeta(r)}{\zeta(2(r+1))}E_{r}(1)\sum_{i=0}^{r}\binom{r}{i}\gamma^{-(i+j+1)}\mathrm{Li}_{i+2}(-u)\Gamma(i+j+1)+O(\gamma^{-j-r-3/2})$
$\leq\left(\frac{\zeta(r+1)K_{r}(1)}{\zeta(r+2)}+\frac{\zeta^{2}(r+1)\zeta(r+2)}{\zeta(2(r+1))}\zeta(r)E_{r}(1)\right)\sum_{i=0}^{r}\binom{r}{i}\gamma^{-(i+j+1)}\mathrm{Li}_{i+2}(-u)\Gamma(i+j+1)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+O(\gamma^{-j-r-3/2}).$ (4.9)
Thus, by combining equations (4.1) and (4.1), we deduce that (4.1) becomes
$\frac{\partial^{j}}{\partial\gamma^{j}}F(\gamma,u)=-\frac{\partial^{j}}{\partial\gamma^{j}}\mathcal{F}_{1}(\gamma,u)+\frac{\partial^{j}}{\partial\gamma^{j}}\mathcal{F}_{2}(\gamma,u)=(-1)^{j+1}\mathfrak{h}_{1,j}(\gamma,u)+(-1)^{j}\mathfrak{h}_{2,j}(\gamma,u)$
$\leq$ $\bigg|\left(\frac{\zeta(r+1)K_{r}(1)}{\zeta(r+2)}+\frac{\zeta^{2}(r+1)\zeta(r+2)}{\zeta(2(r+1))}\zeta(r)E_{r}(1)\right)\sum_{i=0}^{r}\binom{r}{i}\gamma^{-(i+j+1)}\mathrm{Li}_{i+2}(-u)\Gamma(i+j+1)$
$-\zeta(r+1)\mathrm{Li}_{r+2}(-u)\Gamma(r+j+1)\gamma^{-(r+j+1)}\bigg|+O(\gamma^{-j-r-3/2})$
$\leq$ $N(r)\mathrm{Li}_{r+2}(-u)\Gamma(r+j+1)\gamma^{-(r+j+1)}+O(\gamma^{-j-r-3/2}),$ (4.10)
where
$N(r)\coloneqq\left|\frac{\zeta(r+1)K_{r}(1)}{\zeta(r+2)}+\frac{\zeta^{2}(r+1)\zeta(r+2)}{\zeta(2(r+1))}\zeta(r)E_{r}(1)-\zeta(r+1)\right|.$ (4.11)
It is important to note that due to the oscillatory behaviour of the gap function $\Delta_{f}(n)$ for $f(n)=\sigma_{r}(n)$, together with the fact that we can only derive an upper bound for $D_{2}(s,r)$, we obtain an average order for the derivative of $F(\gamma,u)$ rather than an exact formula.
Thus, for the saddle-point $n$, we have
$n=-F_{\gamma}(\tau,u)=$ $O\left(N(r)\mathrm{Li}_{r+2}(-u)\Gamma(r+2)\tau^{-(r+2)}\right)$ (4.12)
whereas
$B^{2}=F_{\gamma\gamma}(\tau,u)=$ $O\left(N(r)\mathrm{Li}_{r+2}(-u)\Gamma(r+3)\tau^{-(r+3)}\right).$
For the third derivative appearing in the error term of (4.3), we have
$\frac{\partial^{3}}{\partial\theta^{3}}F(\tau+i\theta,u)=-i\sum_{k=1}^{\infty}\frac{k^{3}\Delta_{f}(k)(1-u^{-1}e^{k(\tau+i\theta)})}{(1+u^{-1}e^{k(\tau+i\theta)})^{3}}.$
The following estimation technique is adapted from [23]. Let $k_{0}=\tau^{-(1+c)}$ for some $c>0$, and for simplicity of notation, set $v=u^{-1}$. We consider two cases: first, when $k$ is large i.e., $k>k_{0}$; and second, when $k\leq k_{0}$. Beginning with the large sum, we obtain
$\left|-i\sum_{k>k_{0}}\frac{k^{3}\Delta_{f}(k)(1-ve^{k(\tau+i\theta)})}{(1+ve^{k(\tau+i\theta)})^{3}}\right|\leq\sum_{k>k_{0}}\frac{k^{3}|\Delta_{f}(k)|(1+ve^{k\tau})}{|1+ve^{k(\tau+i\theta)}|^{3}}$
$\leq$ $\sum_{k>k_{0}}\frac{k^{3}|\Delta_{f}(k)|(1+ve^{k\tau})}{(ve^{k\tau}-1)^{3}}=O\left(\sum_{k>k_{0}}\frac{k^{3}|\Delta_{f}(k)|}{e^{k\tau}}\right)=O\left(\tau^{-r-4}\right),$ (4.13)
where we have applied the upper bound (see, for instance, [28])
$\sigma_{r}(n)<n^{r}(1+2^{-r}+3^{-r}+\cdots)<n^{r}\zeta(r).$
For the remaining case, we observe that
$\left|1+v^{k(\tau+i\theta)}\right|\geq(1+ve^{k\tau})\cos{\left(\frac{k\theta}{2}\right)}.$
Thus, we obtain
$\left|-i\sum_{k\leq k_{0}}\frac{k^{3}\Delta_{f}(k)(1-u^{-1}e^{k(\tau+i\theta)})}{(1+u^{-1}e^{k(\tau+i\theta)})^{3}}\right|\leq\sum_{k\leq k_{0}}\frac{k^{3}|\Delta_{f}(k)|(1+ve^{k\tau})}{|1+ve^{k(\tau+i\theta)}|^{3}}$
$\leq$ $\sum_{k\leq k_{0}}\frac{k^{3}|\Delta_{f}(k)|(1+ve^{k\tau})}{(1+ve^{k(\tau+i\theta)})^{3}}\left(1+O((k\theta)^{2})\right)\leq\sum_{k\leq k_{0}}\frac{k^{3}|\Delta_{f}(k)|}{ve^{k\tau}}\left(1+O((k\theta)^{2})\right)$
$\leq$ $\sum_{k\geq 1}\frac{k^{3}|\Delta_{f}(k)|}{ve^{k\tau}}+O\left(\sum_{k\geq 1}\frac{k^{5}|\Delta_{f}(k)|\theta^{2}}{e^{k\tau}}\right).$
Applying the Mellin transform and converse mapping, we have
$\left|-i\sum_{k=1}^{\infty}\frac{k^{3}\Delta_{f}(k)(1-u^{-1}e^{k(\tau+i\theta)})}{(1+u^{-1}e^{k(\tau+i\theta)})^{3}}\right|=O\left(\tau^{-r-4}+\tau^{-r-6}\theta^{2}\right).$
Combining this with (4.1), we get
$\frac{\partial^{3}}{\partial\theta^{3}}F(\tau+i\theta,u)=O(\tau^{-r-4}),$
for $|\theta|\leq\tau^{1+3r/7}$. All these together leads to the expansion
$F(\tau+i\theta,u)=F(\tau,u)+in\theta-\frac{B^{2}}{2}\theta^{2}+O(\tau^{-r-4}\theta^{3}).$
Thus for the integral $\Omega_{1,f}$, we obtain
$\frac{e^{n\tau}}{2\pi}\int_{-\theta_{n}}^{\theta_{n}}\exp(in\theta+F(\tau+i\theta,u))d\theta$
$=$ $\frac{e^{n\tau+F(\tau,u)}}{2\pi}\int_{-\theta_{n}}^{\theta_{n}}\exp\left(-\frac{B^{2}}{2}\theta^{2}+O(\tau^{2r/7-1})\right)d\theta$
$=$ $\frac{e^{n\tau+F(\tau,u)}}{2\pi}\left(\int_{-\theta_{n}}^{\theta_{n}}\exp\left(-\frac{B^{2}}{2}\theta^{2}\right)d\theta\right)\left(1+O(\tau^{2r/7-1})\right).$
Finally, we transform the integral into a Gaussian integral and arrive at
$\int_{-\theta_{n}}^{\theta_{n}}\exp\left(-\frac{B^{2}}{2}\theta^{2}\right)d\theta=$ $\int_{-\infty}^{\infty}\exp\left(-\frac{B^{2}}{2}\theta^{2}\right)d\theta-2\int_{\theta_{n}}^{\infty}\exp\left(-\frac{B^{2}}{2}\theta^{2}\right)d\theta$
$=$ $\frac{\sqrt{2\pi}}{B}+O\left(\int_{\theta_{n}}^{\infty}\exp\left(-\frac{B^{2}\tau^{1+3r/7}}{2}\theta\right)d\theta\right)$
$=$ $\frac{\sqrt{2\pi}}{B}+O\left(\tau^{-1-3r/7}B^{-2}\exp\left(-\frac{\tau^{-r/7-1}}{2}\right)\right),$
completing the proof. ∎
4.2. Estimate of $\Omega_{2,f}$
We proceed to prove the following lemma, which establishes an upper bound for the integrand in (4.2) when $f(n)=\sigma_{r}(n)$ for fixed $r\geq 2$.
Lemma 4.2.
For a fixed $r\geq 2$ with $f(n)=\sigma_{r}(n)$, the error term is bounded as follows
$|\Omega_{2,\sigma_{r}}|=O\left(\exp\left(n\tau+F(\tau,u)-c_{r}\tau^{-r/7}\right)\right),$
where $c_{r}$ is a constant uniformly in $u$.
We first note the elementary inequality $\sigma_{r}(n)>n^{r}$ for $n\geq 2$ and $r\geq 1$, which simplifies the analysis in this section. Additionally, we incorporate the following lemma by Li and Chen [21, Lemma 2.5] as a key element of the preceding argument.
Lemma 4.3 (Li-Chen, [21]).
Let $k\geq 1$ be fixed. Then for $\xi>0$, we have
$\sum_{n\geq 1}n^{k-1}e^{-n\xi}(1-\cos(ny))\geq\varrho\left(\frac{e^{-\xi}}{(1-e^{-\xi})^{k}}-\frac{e^{-\xi}}{\left|1-e^{-\xi-iy}\right|^{k}}\right),$
where $\varrho$ is a positive constant which depends only on $k$.
We are now ready to prove the following result, which will directly imply Lemma 4.2. In this context, it is noteworthy that, owing to the inequality $\sigma_{r}(n)>n^{r}$, for $n\geq 2$ and $r\geq 1$, the following lemma bears resemblance to Lemma 7 in [22].
Lemma 4.4.
For a fixed $r\geq 2$ and any real $y$ satisfying $\theta_{0}\leq|y|\leq\pi$, we have
$\frac{|Q_{\sigma_{r}}(e^{\tau+iy},u)|}{Q_{\sigma_{r}}(e^{-\tau,u})}\leq\left(\exp\left(-\frac{2u(r-1)}{(1+u)^{2}}\varrho\left(\frac{1}{4}\tau^{-{r}/{7}}\right)\right)\right),$
where $\varrho$ is a constant depends on $r$.
Proof.
Utilising the inequality $\sigma_{r}(n)>n^{r}$ for $n\geq 2$ and $r\geq 2$, we can simplify
$|\Delta_{f}(n)|=|\sigma_{r}(n+1)-\sigma_{r}(n)|\geq c((n+1)^{r}-n^{r})=c\sum_{i=0}^{r-1}\binom{r}{i}n^{i},$
for some constant $c\geq 0$. However, for $c=0$ in the above inequality, the following argument becomes trivial. So, we only consider the case where $c>0$.
By the mean value theorem, there exists a $\xi\in(n,n+1)$ such that $(n+1)^{r}-n^{r}=r\xi^{r-1}$. This implies (see [21, Lemma 2.6])
$|\Delta_{f}(n)|\geq r\xi^{r-1}-1=r\xi^{r}+\xi^{r}-1=(r-1)\xi^{r-1}+\xi^{r-1}-1>(r-1)n^{r-1}.$ (4.14)
Next, we express
$\left|1+ue^{-n(\tau+iy)}\right|^{2}=(1+ue^{-n\tau-niy})(1-ue^{-n\tau+niy})=(1+ue^{n\tau})^{2}-2ue^{-n\tau}(1-\cos(ny)).$
Using the above identities, we obtain
$\left(\frac{|Q_{\sigma_{r}}(e^{\tau+iy},u)|}{Q_{\sigma_{r}}(e^{-\tau,u})}\right)^{2}=$ $\prod_{n=1}^{\infty}\left(1-\frac{2ue^{-n\tau}(1-\cos(ny))}{(1+ue^{-n\tau})^{2}}\right)^{\Delta_{f}(n)}$
$\leq$ $\exp\left(\sum_{n=1}^{\infty}|\Delta_{f}(n)|\log\left|1-\frac{2ue^{-n\tau}(1-\cos(ny))}{(1+ue^{-n\tau})^{2}}\right|\right)$
$\leq$ $\exp\left(-\sum_{n=1}^{\infty}|\Delta_{f}(n)|\left|\frac{2ue^{-n\tau}(1-\cos(ny))}{(1+ue^{-n\tau})^{2}}\right|\right)$
$\leq$ $\exp\left(-\frac{2u}{(1+u)^{2}}\sum_{n=1}^{\infty}|\Delta_{f}(n)|e^{-n\tau}(1-\cos(ny))\right)$
$\leq$ $\exp\left(-\frac{2u(r-1)}{(1+u)^{2}}\sum_{n=1}^{\infty}n^{r-1}e^{-n\tau}(1-\cos(ny))\right).$
Finally, applying Lemma 4.3, yields
$\left(\frac{|Q_{\sigma_{r}}(e^{\tau+iy},u)|}{Q_{\sigma_{r}}(e^{-\tau,u})}\right)^{2}\leq\exp\left(-\frac{2u(r-1)}{(1+u)^{2}}\varrho\left(\frac{e^{-\tau}}{(1-e^{-\tau})^{r}}-\frac{e^{-\tau}}{\left|1-e^{-\tau-iy}\right|^{r}}\right)\right).$ (4.15)
Beginning with a few standard trigonometric identities from [21], we express
$\left|1-e^{-\tau-i\theta}\right|^{r}=$ $((1-e^{-\tau})^{2}+2e^{-\tau}(1-\cos\theta))^{\tfrac{r}{2}}$
$\geq$ $(1-e^{-\tau})^{r}\left(1+\frac{2e^{-\tau}}{(1-e^{-\tau})^{2}}(1-\cos\theta_{n})\right)^{\tfrac{r}{2}},$ (4.16)
where $\theta_{n}\leq|\theta|\leq\pi$. By employing the Taylor series expansion of $\cos{\theta}$, we obtain
$1-\cos\theta_{n}=\frac{1}{2}\theta_{n}^{2}+O(\theta_{n}^{4})=\frac{1}{2}\tau^{2+{6r}/{7}}+O(\theta_{n}^{4+12r/7}),$
and similarly, $e^{-\tau}=1-\tau+O(\tau^{2})$. Consequently, we derive
$\frac{2e^{-\tau}}{(1-e^{-\tau})^{2}}(1-\cos\theta_{n})=\frac{2(1+O(\tau))(\tfrac{1}{2}\tau^{2+{6r}/{7}}+O(\theta_{n}^{4+12r/7}))}{\tau^{2}(1+O(\tau))}=\tau^{{6r}/{7}}(1+O(\tau)).$
Substituting the above expression into (4.2) gives
$\left|1-e^{-\tau-i\theta}\right|^{r}\geq(1-e^{-\tau})^{r}\left(1+\tau^{{6r}/{7}}(1+O(\tau))\right)^{\tfrac{r}{2}}$
$=(1-e^{-\tau})^{r}\left(1+\frac{r\tau^{{6r}/{7}}}{2}+O\left(\tau^{{6r}/{7}+1}+\tau^{{12r}/{7}}\right)\right).$
Therefore, the inner sum of (4.15) simplifies to
$\frac{e^{-\tau}}{(1-e^{-\tau})^{r}}-\frac{e^{-\tau}}{\left|1-e^{-\tau-iy}\right|^{r}}$ $\geq\frac{e^{-\tau}}{(1-e^{-\tau})^{r}}\left(1-\frac{1}{1+\frac{r\tau^{{6r}/{7}}}{2}+O\left(\tau^{{6r}/{7}+1}+\tau^{{12r}/{7}}\right)}\right)$
$=\frac{e^{-\tau}}{(1-e^{-\tau})^{r}}\left(\frac{r\tau^{{6r}/{7}}}{2}+O\left(\tau^{{6r}/{7}+1}+\tau^{{12r}/{7}}\right)\right)$
$=\frac{1-O(\tau)}{\tau^{r}(1-O(\tau))}\left(\frac{r\tau^{{6r}/{7}}}{2}+O\left(\tau^{{6r}/{7}+1}+\tau^{{12r}/{7}}\right)\right)$
$\geq\frac{r\tau^{-{r}/{7}}}{2}+O\left(\tau^{1-{r}/{7}}+\tau^{12r/7}\right)\gg\frac{1}{4}\tau^{-\tfrac{r}{7}}.$
Substituting this bound into (4.15), we conclude the proof. ∎
Proof of Lemma 4.2. By the definition given in (4.2) for $f(n)=\sigma_{r}(n)$, the following inequity holds
$|\Omega_{2,\sigma_{r}}|\leq\frac{e^{n\tau}}{2\pi}\int_{\theta_{n}}^{\pi}\left|{Q}_{\sigma_{r}}(e^{-\tau-i\theta},u)\right|d\theta=\frac{e^{n\tau+F(\tau,u)}}{\pi}\int_{\theta_{n}}^{\pi}\frac{|Q_{\sigma_{r}}(e^{-\tau-i\theta},u)|}{Q_{\sigma_{r}}(e^{-\tau},u)}d\theta.$
By employing Lemma 4.4, we arrive at
$|\Omega_{2,\sigma_{r}}|\ll\exp\left(n\tau+F(\tau,u)-c_{r}\tau^{-r/7}\right),$
where $c_{r}$ is a constant that remains uniform in $u$, as noted. ∎
5. The Moment Generating Function
In this section, we employ the estimates of $\Omega_{1,f}$ and $\Omega_{2,f}$ to derive
$\mathcal{Q}_{f,n}(u)=\frac{e^{n\tau+F(\tau,u)}}{\sqrt{2\pi B}}\left(1+O\left(\tau^{2r/7-1}\right)\right),$
where $r\geq 2$ is fixed and $B^{2}=F_{\gamma\gamma}(\tau,u)$.
We adopt the method outlined in [22, §4.3], by considering the asymptotic expansion of the moment generating function associated with the random variable ${\varpi}_{f,n}$. Let $M_{f,n}(\theta)=\mathbb{E}(e^{(\varpi_{f,n}-\mu_{f,n})\theta/\nu_{f,n}})$ denote the moment generating function for real $\theta$, and let $\mu_{f,n}$ and $\nu_{f,n}$ represent the mean and standard deviation, as defined in (2.4) and (2.5), respectively. This leads to the following result.
Lemma 5.1.
Let $r\geq 2$ be fixed and $f(n)=\sigma_{r}(n)$ be the generalised divisor function. Then as $n\to\infty$, for bounded $\theta$, the following holds
$M_{f,n}(\theta)=\exp\left(\frac{\theta^{2}}{2}+O\left(n^{-2r/7(r+2)+1/(r+2)}\right)\right).$
Proof.
We begin by noting that
$M_{f,n}(\theta)=$ $\exp\left(-\frac{\mu_{f,n}\theta}{\nu_{f,n}}\right)\frac{\mathcal{Q}_{f,n}(e^{\theta/\nu_{f,n}})}{\mathcal{Q}_{f,n}(1)}$
$=$ $\sqrt{\frac{B^{2}(\tau(n,1),1)}{B^{2}(\tau(n,e^{\theta/\nu_{f,n}}),e^{\theta/\nu_{f,n}})}}\exp\Bigg(-\frac{\mu_{f,n}\theta}{\nu_{f,n}}+n\tau(n,e^{\theta/\nu_{f,n}})$
$\quad\quad\quad+{F}(\tau(n,e^{\theta/\nu_{f,n}}),e^{\theta/\nu_{f,n}})-n\tau(n,1)-{F}(\tau(n,1),1)+O(\tau^{2r/7-1})\Bigg).$ (5.1)
For simplicity of notation, we define $\tau=\tau(n,u)$ which is a function of $n$ and $u$ with respect to $u$. By employing the saddle-point solution given in (4.12), we obtain (following the approach of [23])
$\tau_{u}=\tau_{u}(n,u)=-\frac{{F}_{u\gamma}(\tau,u)}{{F}_{\gamma\gamma}(\tau,u)}=\frac{\sum_{k=1}^{\infty}\frac{k\Delta_{f}(k)e^{k\tau}}{(e^{k\tau+u})^{2}}}{\sum_{k=1}^{\infty}\frac{uk^{2}\Delta_{f}(k)e^{kr}}{(e^{kr}+u)^{2}}},$
and similarly,
$\tau_{uu}=\tau_{uu}(n,u)=\frac{1}{{F}_{\gamma\gamma}(\tau,u)^{3}}\big(-{F}_{\gamma\gamma\gamma}(\tau,u){F}_{u\gamma}(\tau,u)^{2}$ $+2{F}_{u\gamma\gamma}(\tau,u){F}_{u\gamma}(\tau,u){F}_{\gamma\gamma}(\tau,u)$
$-{F}_{uu\gamma}(\tau,u){F}_{\gamma\gamma}(\tau,u)^{2}\big),$
and
$\tau_{uuu}=\tau_{uuu}(n,u)=$ $\frac{1}{{F}_{\gamma\gamma}(\tau,u)^{5}}\big(-{F}_{uuu\gamma}(\tau,u){F}_{\gamma\gamma}(\tau,u)^{4}$
$+(3{F}_{uu\gamma\gamma}(\tau,u){F}_{u\gamma}(\tau,u)+3{F}_{uu\gamma}(\tau,u){F}_{u\gamma\gamma}(\tau,u)){F}_{\gamma\gamma}(\tau,u)^{3}$
$+\big(-3{F}_{u\gamma\gamma\gamma}(\tau,u){F}_{u\gamma}(\tau,u)^{2}-6{F}_{u\gamma\gamma}(\tau,u)^{2}{F}_{u\gamma}(\tau,u)$
$-3{F}_{uu\gamma}(\tau,u){F}_{\gamma\gamma\gamma}(\tau,u){F}_{u\gamma}(\tau,u)\big)F_{\gamma\gamma}(\tau,u)^{2}$
$+\big({F}_{\gamma\gamma\gamma\gamma}(\tau,u){F}_{u\gamma}(\tau,u)^{3}+9{F}_{u\gamma\gamma}(\tau,u){F}_{\gamma\gamma\gamma}(\tau,u){F}_{u\gamma}(\tau,u)^{2}\big){F}_{\gamma\gamma}(\tau,u)$
$-3{F}_{u\gamma}(\tau,u)^{3}{F}_{\gamma\gamma\gamma}(\tau,u)^{2}\big).$
As it is evident from the preceding expression, it is necessary to estimate the partial derivative of ${F}_{\gamma}(\tau,u)$ with respect to $u$. To achieve this, we apply the estimate established in Lemma 4.1 and proceed using a method analogous to that employed previously with the converse mapping. We provide a detailed illustration for ${F}_{u\gamma}(\tau,u)$; the remaining cases follow in a similar manner. For $j=1$, from (4.1) we write
${F}_{\gamma}(\gamma,u)=\mathfrak{h}_{1,1}(\gamma,u)-\mathfrak{h}_{2,1}(\gamma,u)=\sum_{n=1}^{\infty}\sigma_{r}(n+1)n\sum_{\ell=1}^{\infty}(-u)^{\ell}e^{-n\ell\gamma}-\sum_{n=1}^{\infty}\sigma_{r}(n)n\sum_{\ell=1}^{\infty}(-u)^{\ell}e^{-n\ell\gamma}.$
Then the Mellin transform of $\mathfrak{h}_{1,1}(\gamma,u)-\mathfrak{h}_{2,1}(\gamma,u)$ is given by
$\mathcal{H}_{1,1}(s,u)-\mathcal{H}_{2,1}(s,u)=O\Bigg(\zeta(r+1)\mathrm{Li}_{s}(-u)\Gamma(s)\bigg(\sum_{i=0}^{r}\binom{r}{i}\frac{\zeta(s-i-1)}{\zeta(s-i+r)}K_{r}(s-i-1)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\sum_{i=0}^{r}\binom{r}{i}\frac{\zeta(\sigma-i-1)\zeta(\sigma-i-1+r)\zeta(\sigma-i+r)}{\zeta(2(\sigma-i-1+r))}\zeta(r)E_{r}(\sigma-i-1)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\zeta(s-1)\zeta(s-1-r)}{\zeta(r+1)}\bigg)\Bigg),$
where $s=\sigma+it$ with $\sigma,t\in\mathbb{R}$, and $K_{r}(s)$ and $E_{r}(s)$ are defined in (3.9) and (3.17), respectively. Differentiating with respect to $u$, we obtain
$\frac{\partial}{\partial u}(\mathcal{H}_{1,1}(s,u)-\mathcal{H}_{2,1}(s,u))=O\Bigg(\frac{1}{u}\mathrm{Li}_{s-1}(-u)\Gamma(s)\zeta(r+1)\bigg(\sum_{i=0}^{r}\binom{r}{i}\frac{\zeta(s-i-1)}{\zeta(s-i+r)}K_{r}(s-i-1)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\sum_{i=0}^{r}\binom{r}{i}\frac{\zeta(\sigma-i-1)\zeta(\sigma-i-1+r)\zeta(\sigma-i+r)}{\zeta(2(\sigma-i-1+r))}\zeta(r)E_{r}(\sigma-i-1)$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{\zeta(s-1)\zeta(s-1-r)}{\zeta(r+1)}\bigg)\Bigg).$
By applying Theorem 2.3, yields
$F_{\gamma u}(\tau,u)=$ $O\Bigg(\frac{1}{u}\zeta(r+1)\bigg(\sum_{i=0}^{r}\binom{r}{i}\mathrm{Li}_{i+1}(-u)\Gamma(i+2)\frac{K_{r}(1)}{\zeta(r+2)}\tau^{-(i+2)}$
$+$ $\sum_{i=0}^{r}\binom{r}{i}\mathrm{Li}_{i+1}(-u)\Gamma(i+2)\frac{\zeta(r+1)\zeta(r+2)\zeta(r)E_{r}(1)}{\zeta(2(r+1))}\tau^{-(i+2)}-\mathrm{Li}_{r+1}(-u)\Gamma(r+2)\tau^{-(r+2)}\bigg)\Bigg)$
$=$ $O\left(\frac{N(r)}{u}\mathrm{Li}_{r+1}(-u)\tau^{-(r+2)}\Gamma(r+2)\right)=O(n),$
where $N(r)$ is defined in (4.11). In a similar fashion, we derive the remaining partial derivatives and get
$F_{u}(\tau,u)=\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)}{e^{k\tau}+u}=O\left(\frac{N(r)}{u}\mathrm{Li}_{r+1}(-u)\tau^{-(r+1)}\Gamma(r+1)\right)=O(n^{\frac{r+1}{r+2}}),$
$F_{uu}(\tau,u)=-\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)}{(e^{k\tau}+u)^{2}}=O\left(\frac{N(r)}{u^{2}}(\mathrm{Li}_{r}(-u)-\mathrm{Li}_{r+1}(-u))\Gamma(r+1)\tau^{-(r+1)}\right)=O(n^{\frac{r+1}{r+2}}),$
$F_{\gamma\gamma}(\tau,u)=\sum_{k=1}^{\infty}\frac{uk^{2}\Delta_{f}(k)e^{k\tau}}{(e^{k\tau}+u)^{2}}=O\left(N(r)\mathrm{Li}_{r+2}(-u)\Gamma(r+3)\tau^{-(r+3)}\right)=O(n^{\frac{r+3}{r+2}}),$
$F_{uuu}(\tau,u)=\sum_{k=1}^{\infty}\frac{2\Delta_{f}(k)}{(e^{k\tau}+u)^{3}}=O(\tau^{-(r+1)})=O(n^{\frac{r+1}{r+2}}),$
$F_{uu\gamma}(\tau,u)=\sum_{k=1}^{\infty}\frac{2k\Delta_{f}(k)e^{k\tau}}{(e^{k\tau}+u)^{3}}=O(\tau^{-(r+2)})=O(n),$
$F_{u\gamma\gamma}(\tau,u)=\sum_{k=1}^{\infty}\frac{k^{2}\Delta_{f}(k)e^{k\tau}(e^{k\tau}-u)}{(e^{k\tau}+u)^{3}}=O(\tau^{-(r+3)})=O(n^{\frac{r+3}{r+2}}),$
$F_{\gamma\gamma\gamma}(\tau,u)=-\sum_{k=1}^{\infty}\frac{uk^{3}\Delta_{f}(k)e^{k\tau}(e^{k\tau}-u)}{(e^{k\tau}+u)^{3}}=O(\tau^{-(r+4)})=O(n^{\frac{r+4}{r+2}}),$
$F_{uuu\gamma}(\tau,u)=-\sum_{k=1}^{\infty}\frac{6k\Delta_{f}(k)e^{k\tau}}{(e^{k\tau}+u)^{4}}=O(\tau^{-(r+2)})=O(n),$
$F_{uu\gamma\gamma}(\tau,u)=-\sum_{k=1}^{\infty}\frac{2k^{2}\Delta_{f}(k)e^{k\tau}(e^{k\tau}-u)}{(e^{k\tau}+u)^{4}}=O(\tau^{-(r+3)})=O(n^{\frac{r+3}{r+2}}),$
$F_{u\gamma\gamma\gamma}(\tau,u)=-\sum_{k=1}^{\infty}\frac{k^{3}\Delta_{f}(k)e^{k\tau}(e^{2k\tau}-4ue^{k\tau}+u^{2})}{(e^{k\tau}+u)^{4}}=O(\tau^{-(r+4)})=O(n^{\frac{r+4}{r+2}}),$
$F_{\gamma\gamma\gamma\gamma}(\tau,u)=\sum_{k=1}^{\infty}\frac{uk^{4}\Delta_{f}(k)e^{k\tau}(e^{2k\tau}-4ue^{k\tau}+u^{2})}{(e^{k\tau}+u)^{4}}=O(\tau^{-(r+5)})=O(n^{\frac{r+5}{r+2}}).$
The above bounds imply that the derivatives $\tau_{u}$, $\tau_{uu}$, and $\tau_{uuu}$ are each $O(n^{-1/(r+2)})$, uniformly with respect to $u$. Expanding $\tau(n,e^{\theta/\nu_{f,n}})$ and $F(\tau(e^{\theta/\nu_{f,n}},n),e^{\theta/\nu_{f,n}})$ around $\theta=0$ yields
$\tau(n,e^{\theta/\nu_{f,n}})-\tau(n,1)=\tau_{u}(n,1)\frac{\theta}{\nu_{f,n}}+\frac{\tau_{u}(n,1)+\tau_{uu}(n,1)}{2}\left(\frac{\theta}{\nu_{f,n}}\right)^{2}+O\left(n^{-1/(r+2)}\frac{|\theta|^{3}}{\nu^{3}_{f,n}}\right)$
and
$F(\tau(e^{\theta/\nu_{f,n}},n),e^{\theta/\nu_{f,n}})-F((\tau,1),1)$
$=$ $F_{\gamma}(\tau(n,1),1)\tau_{u}(n,1)+F_{u}(\tau(n,1),1)\frac{\theta}{\nu_{f,n}}+\frac{1}{2}\bigg(F_{\gamma}(\tau(n,1),1)(\tau_{u}(n,1)+\tau_{uu}(n,1))$
$+F_{\gamma\gamma}(\tau(n,1),1)\tau_{u}(n,1)^{2}+2F_{u\gamma}(\tau(n,1),1)\tau_{u}(n,1)+F_{u}(\tau(n,1),1)$
$+F_{uu}(\tau(n,1),1)\bigg)\left(\frac{\theta}{\nu_{f,n}}\right)^{2}+O\left(n^{\frac{r+1}{r+2}}\frac{|\theta|^{3}}{\nu^{3}_{f,n}}\right),$
respectively. Substituting these expansions into the exponential in (5), we obtain
$\left(n\tau_{u}(n,1)+F_{\gamma}(\eta,1)\tau_{u}(n,1)+F_{u}(\eta,1)-\mu_{f,n}\right)\frac{\theta}{\nu_{f,n}}+\frac{1}{2}\bigg(n(\tau_{u}(n,1)+\tau_{uu}(n,1))+F_{\gamma}(\eta,1)(\tau_{u}(n,1)$
$+\tau_{uu}(n,1))+F_{\gamma\gamma}(\eta,1)\tau_{u}(n,1)^{2}+2F_{u\gamma}(\eta,1)\tau_{u}(n,1)+F_{u}(\eta,1)+F_{uu}(\eta,1)\bigg)\left(\frac{\theta}{\nu_{f,n}}\right)^{2}$
$+O\left(n^{\frac{r+1}{r+2}}\frac{|\theta^{3}|}{\nu^{3}_{f,n}}+n^{-2r/7(r+2)+1/(r+2)}\right).$
Here, we have written $\eta=\tau(n,1)$ for brevity. Recalling that $n=-F_{\gamma}(\eta,1)$ and $\tau_{u}(n,1)=-\dfrac{F_{u\gamma}(\eta,1)}{F_{\gamma\gamma}(\eta,1)}$, we may simplify the final expression to derive
$(F_{u}(\eta,1)-\mu_{f,n})\frac{\theta}{\nu_{f,n}}+\frac{1}{2}\bigg(F_{u}(\eta,1)+F_{uu}(\eta,1)-\frac{F_{u\gamma}(\eta,1)^{2}}{F_{\gamma\gamma}(\eta,1)}\bigg)\left(\frac{\theta}{\nu_{f,n}}\right)^{2}$
$+O\left(n^{\frac{r+1}{r+2}}\frac{|\theta^{3}|}{\nu^{3}_{f,n}}+n^{-2r/7(r+2)+1/(r+2)}\right).$
In a similar manner, we obtain that
$\frac{B^{2}(\tau(n,1),1)}{B^{2}(\tau(n,e^{\theta/\nu_{f,n}}),e^{\theta/\nu_{f,n}})}=1+O\left(\frac{|\theta|}{\nu_{f,n}}\right).$
Thus, we derive the following asymptotic expression for the moment generating function in (5):
$M_{f,n}(\theta)=$ $\exp\Bigg((F_{u}(\eta,1)-\mu_{f,n})\frac{\theta}{\nu_{f,n}}+\frac{1}{2}\bigg(F_{u}(\eta,1)+F_{uu}(\eta,1)-\frac{F_{u\gamma}(\eta,1)^{2}}{F_{\gamma\gamma}(\eta,1)}\bigg)\left(\frac{\theta}{\nu_{f,n}}\right)^{2}$
$+O\left(n^{\frac{r+1}{r+2}}\frac{|\theta^{3}|}{\nu^{3}_{f,n}}+n^{-2r/7(r+2)+1/(r+2)}\right)\Bigg).$
Recall that we choose $\mu_{f,n}$ and $\nu_{f,n}$ in (2.4) and (2.5) such that
$\mu_{f,n}=F_{u}(\eta,1)=\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)}{e^{\eta k}+1}\quad\text{and}\quad$
$\nu^{2}_{f,n}=F_{u}(\eta,1)+F_{uu}(\eta,1)-\frac{F_{u\gamma}(\eta,1)^{2}}{F_{\gamma\gamma}(\eta,1)}=\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)e^{\eta k}}{(e^{\eta k}+1)^{2}}-\frac{\left(\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)ke^{\eta k}}{(e^{\eta k}+1)^{2}}\right)^{2}}{\sum_{k=1}^{\infty}\frac{\Delta_{f}(k)k^{2}e^{\eta k}}{(e^{\eta k}+1)^{2}}}.$ (5.2)
From the definitions of $\mu_{f,n}$ and $\nu^{2}_{f,n}$ in (2.4) and (2.5) respectively, we deduce that
$M_{f,n}(\theta)=$ $\exp\left(\frac{\theta^{2}}{2}+O\left(n^{-(r+1)/(r+2)}|\theta|+n^{-(r+1)/(r+2)}|\theta|^{3}+n^{-2r/7(r+2)+1/(r+2)}\right)\right)$
$=$ $\exp\left(\frac{\theta^{2}}{2}+O\left(n^{-2r/7(r+2)+1/(r+2)}\right)\right),$
for bounded $\theta$. ∎
5.1. Formulae for the mean and variance
By applying Lemma 5.1 and Curtiss’ theorem [7], we establish that the distribution of ${\varpi}_{f,n}$ is indeed asymptotically normal. In this section, we derive an upper bound for its mean and variance, providing an explicit formula for the constant as a function of $r$.
From (5) we write
$\mu_{f,n}=F_{u}(\eta,1)=O\left(N(r)\mathrm{Li}_{r+1}(-1)\Gamma(r+1)/\eta^{(r+1)}\right)$
and
$\nu^{2}_{f,n}=$ $F_{u}(\eta,1)+F_{uu}(\eta,1)-\frac{F_{u\gamma}(\eta,1)^{2}}{F_{\gamma\gamma}(\eta,1)}$
$=$ $O\left(\left(N(r)\mathrm{Li}_{r}(-1)\Gamma(r+1)-N(r)\frac{\mathrm{Li}^{2}_{r}(-1)\Gamma^{2}(r+2)}{\mathrm{Li}_{r+1}(-1)\Gamma(r+3)}\right)\frac{1}{\eta^{(r+1)}}\right).$
Utilising the saddle-point solution in (4.12), we write
$\frac{1}{\eta}=\frac{1}{\tau(n,1)}=O\left(\frac{n^{\frac{1}{r+2}}}{{N(r)\mathrm{Li}_{r+2}(-1)\Gamma(r+2)}}\right)$
then
$\mu_{f,n}=O\left(C_{\mu}(r)n^{\frac{r+1}{r+2}}\right)$
and
$\nu^{2}_{f,n}=O\left(C_{\nu}(r)n^{\frac{r+1}{r+2}}\right),$
where
$C_{\mu}(r)\sim-N(r){\mathrm{Li}_{r+1}(-1)\Gamma(r+1)}=N(r)(1-2^{r+1})\Gamma(r+1)\zeta(r+1)$ (5.3)
and
$C_{\sigma}(r)\sim$ $-N(r)\left({\mathrm{Li}_{r}(-1)\Gamma(r+1)-\frac{\mathrm{Li}^{2}_{r}(-1)\Gamma^{2}(r+2)}{\mathrm{Li}_{r+1}(-1)\Gamma(r+3)}}\right)$
$=$ $N(r)\zeta(r)(1-2^{-r})\left(\Gamma(r+1)-\frac{\Gamma(r+2)\zeta(r)}{\Gamma(r+3)\zeta(r+1)}\left(1-\frac{1}{2^{r+1}-1}\right)\right),$ (5.4)
$N(r)$ is given in (4.11) and
$-\mathrm{Li}_{r}(-1)=(1-2^{-r})\zeta(r+1).$
In this context, it is worth noting that the key constant of interest is $N(r)$, which emerges from a detailed analysis of the Dirichlet series with the gap function $\mathcal{D}_{f}(\Delta;s)$.
5.1.1. The distribution of the tail
We conclude the proof by bounding the tail of the distribution, following the approach of Lipnik et al. [22], which in turn was adapted from Hwang [16]. For $\theta=o(n^{\frac{r+1}{6(r+2)}})$, we apply the Chernoff bound (see, for instance, [6]) to obtain
$\mathbb{P}\left(\frac{\varpi_{f,n}-\mu_{f,n}}{\sigma_{f,n}}\geq x\right)\leq e^{-\theta x}M_{f,n}(\theta)=e^{-t\theta+\theta^{2}/2}\left(1+O\left(n^{\frac{r+1}{2(r+2)}}(|\theta|+|\theta|^{3})+n^{-2r/7(r+2)+1/(r+2)}\right)\right).$
Setting $T=n^{\frac{r+1}{6(r+2)}}/\log n$ we have two cases. For the first case, let $x\leq T$, and we set $\theta=x$, yielding
$\mathbb{P}\left(\frac{\varpi_{f,n}-\mu_{f,n}}{\sigma_{f,n}}\geq x\right)\leq e^{-x^{2}/2}\left(1+O((\log n)^{-3})\right).$
Similarly, for $x\geq T$, we set $\theta=T$, obtaining
$\mathbb{P}\left(\frac{\varpi_{f,n}-\mu_{f,n}}{\sigma_{f,n}}\geq x\right)\leq e^{-Tx/2}\left(1+O((\log n)^{-3})\right).$
A similar argument can be applied to estimate the probability $\mathbb{P}\left(\frac{\varpi_{f,n}-\mu_{f,n}}{\sigma_{f,n}}\leq-x\right)$.
Acknowledgement. We are grateful to Kyle Pratt and Alexandru Zaharescu for their valuable suggestions and many insightful discussions. A special thanks to Oleksiy Klurman for his comments on an earlier version of this paper. The first author is supported by EPSRC doctoral fellowship #EP/T518049/1.
Rights Retention. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.